### v1

# Feedback-Aware Monte Carlo Tree Search for Efficient Information Seeking in Goal-Oriented Conversations

Abstract
The ability to identify and acquire missing information is a critical component of effective decision making and problem solving. With the rise of conversational artificial intelligence (AI) systems, strategically formulating information-seeking questions becomes crucial and demands efficient methods to guide the search process. We introduce a novel approach to adaptive question-asking through a combination of Large Language Models (LLM) for generating questions that maximize information gain, Monte Carlo Tree Search (MCTS) for constructing and leveraging a decision tree across multiple samples, and a hierarchical feedback mechanism to learn from past interactions. We present two key innovations: (1) an adaptive MCTS algorithm that balances exploration and exploitation for efficient search over potential questions; and (2) a clustering-based feedback algorithm that leverages prior experience to guide future interactions. Each incoming sample is assigned to a cluster based on its semantic similarity with previously observed samples. Our UCT (Upper Confidence bound for Trees) formulation selects optimal questions by combining expected rewards, a function of information gain, with a cluster-specific bonus that decays with depth, to emphasize the importance of early-stage questions that have proven effective for narrowing the solution space in similar samples. Experiments across three domains, including medical diagnosis and troubleshooting, demonstrate that our method leads to an average of 12% improvement in success rates and a 10x reduction in the average number of LLM calls made per conversation for the search process, in comparison to the state of the art. [^1] https://github.com/harshita-chopra/misq-hf
1 Introduction
The field of conversational AI has been witnessing significant advancements in recent years, with a growing focus on developing systems capable of adaptive and efficient information-seeking strategies. As Large Language Models (LLMs) continue to push the boundaries of natural language understanding and generation, researchers are exploring novel approaches to enhance their decision-making capabilities in interactive contexts. One of the core challenges in a conversational system that employs multiple rounds of Q&A to arrive at an answer is to determine the next question to ask. A common solution to this challenge is to construct a contextual decision tree that allows the system to strategically plan which questions to pose and in what order, refining the process to ensure the correct answer or goal is achieved efficiently.
LLMs have emerged as powerful tools for planning using trees Zhao et al. (2024); Putta et al. (2024); Hu et al. (2024). Notably, the Tree of Thoughts (ToT) Yao et al. (2023) approach has demonstrated the potential of leveraging hierarchical structures to improve problem-solving and reasoning in language models. However, expanding the full tree can be computationally inefficient and expensive, especially in domains with large search spaces. One of the fundamental problems lies in designing systems capable of dynamically adapting their interaction strategy to converge on a specific solution efficiently. To address these challenges, Monte Carlo Tree Search (MCTS) has emerged as a promising technique for balancing exploration and exploitation in decision-making processes. MCTS has been successfully applied in various domains, from game-playing AI Silver et al. (2016, 2018) to robotics Dam et al. (2022), and is now being leveraged in natural language processing tasks.
Figure 1: Overview of MISQ-HF. The system combines Monte Carlo Tree Search with hierarchical feedback for efficient information seeking. Green nodes ( ) and red nodes ( ) represent $v^{A}$ and $v^{N}$ respectively, corresponding to possibilities after affirmative and negative answers to a question. The purple arrows indicate the information-seeking question loop (steps 4-8). Towards the end of decision-making, targeted questions are generated, denoted by the loop of dashed arrows (steps 9-6-7-8). Finally, when the user response $a_{t}$ (step 7) confirms the target detection, the conversation is terminated (step 10) and the feedback mechanism updates the reward bonus along the successful questioning trajectory in the decision tree, enabling the system to adapt its question selection strategy for similar future cases.
Building upon this foundation, our work introduces a novel framework that combines MCTS with a hierarchical feedback mechanism to achieve adaptive and efficient information-seeking. Inspired by prior work on uncertainty-aware planning, our approach extends these ideas with three key contributions:
  1. 1.
    MCTS on Decision Tree for Question Generation: Traditional tree-based approaches can struggle with the vast search spaces. By employing MCTS, our system efficiently constructs and explores promising branches of the decision tree without exhaustive expansion. We propose a variant of MCTS that optimizes question selection by considering both immediate and long-term information gain. We use the standard Upper Confidence Bound for Trees (UCT) and propose a modified variant that assigns higher weights to questions that have effectively narrowed the solution space in past interactions.
  2. 2.
    Cluster-Based Feedback Mechanism: We leverage similarity-based clustering based on text embeddings of the initial problem description given by the user. We allow the system to learn from historical successes by propagating cluster-specific bonus rewards through the decision tree after a successful interaction. This bonus reward decays with depth, recognizing that early-stage questions play a crucial role in narrowing down the solution space by targeting high-level information that can effectively eliminate large subsets of possibilities. These questions are often broadly applicable across multiple cases, making them valuable for generalizing strategies. In contrast, later-stage questions tend to be more specific to individual cases, focusing on finer details that depend on the unique context of each interaction.
  3. 3.
    Efficiency through reduced LLM calls: Our method minimizes computational overhead by strategically limiting LLM interactions during tree construction and simulation. We maintain a single decision tree for a given dataset which expands as more data points stream in. We also employ lightweight rollout policies limited by predefined depth during the simulation step of MCTS. The exploration-exploitation balance avoids exhaustive tree expansion and significantly reduces the number of LLM calls without compromising performance.
These innovations collectively enable our system to dynamically adapt its questioning strategy based on historical patterns, achieving greater efficiency and relevance. For instance, the integration of feedback mechanisms ensures that successful strategies for specific clusters are reinforced over time, while depth-aware MCTS ensures computational resources are focused on the most promising paths. An overview of our framework, MISQ-HF (Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback), is illustrated in Figure 1.
We evaluate our approach across diverse conversational tasks. Results demonstrate that our system outperforms existing baselines in both task efficiency and relevance, particularly in scenarios requiring complex reasoning and hierarchical decision-making. Additionally, we also highlight the individual contributions of depth-aware MCTS and cluster-based feedback in enhancing the system’s performance.
2 Related Work
The integration of LLMs with tree search algorithms has proven effective for enhancing multi-step reasoning and decision-making. Guez et al. (2018) introduced MCTSnets, combining tree search with neural networks for tasks like Sokoban. While it focuses on discrete planning, our approach extends MCTS to hierarchical conversations, optimizing question selection with depth-aware strategies and semantic embeddings. Yu et al. (2023) used MCTS for dialogue planning with LLM-based simulations, but their approach differs in using open-loop search without feedback mechanisms and focuses only on persuasion tasks.
Hui and Tu (2024) introduced the Reflection on Search Trees (RoT) framework for generating state-specific guidelines to improve search performance. Unlike RoT, which focuses on reasoning efficiency, our approach maintains a single decision tree while incorporating cluster-specific rewards to guide question selection. Zhu et al. (2021) framed retrieval and answering as a partially observed Markov decision process to adaptively gather evidence, though it does not address learning from feedback in conversational settings. In the clinical domain, Li et al. (2024) proposed MEDIQ, an adaptive framework leveraging LLMs to identify missing information and ask follow-up questions. While MEDIQ emphasizes reliability in high-stakes contexts, our work generalizes by prioritizing early-stage questions to effectively narrow the solution space.
Xie et al. (2024) enhanced LLM reasoning through iterative preference learning inspired by AlphaZero, using MCTS to collect step-level preferences. This aligns with our approach of integrating cluster-specific success bonus rewards into MCTS. Zhang et al. (2024) explored autonomous tree search capabilities in LLMs, demonstrating self-guided reasoning, while our method emphasizes leveraging historical feedback through clustering mechanisms.
Our work draws inspiration from the Uncertainty of Thoughts (UoT) algorithm introduced by Hu et al. (2024), which enhances LLMs’ ability to actively seek information by asking effective questions. UoT combines a simulation approach with uncertainty-based rewards motivated by information gain, and a reward propagation scheme to select optimal questions, demonstrating significant improvements in task performance across various domains and baselines like Chain of Thoughts Wei et al. (2022) and Tree of Thoughts Yao et al. (2023). Building upon these principles, our approach extends the concept by integrating a depth-aware MCTS algorithm that optimizes question selection in hierarchical conversations, and a cluster-based feedback system that leverages historical successes to guide future interactions by adding a depth-aware bonus reward to prioritize early-stage questions. By integrating these components, our approach achieves a dynamic balance between efficiency and adaptability.
3 Methodology
In this section, we present our problem formulation, construction of the decision tree, reward function, our MCTS algorithm, and our hierarchical feedback approach.
3.1 Problem Formulation
The task is modeled as a sequential interaction between a Questioner (an LLM) and an Answerer (a human; here, simulated by an LLM). Let $\Omega$ be the space of possible target items, with an unknown target $\omega\in\Omega$. The interaction proceeds over turns $t=1,\dots,T$, where $T$ is the maximum allowed turns. At each turn, the system asks a question $q_{t}$ and receives an answer $a_{t}$, which is either binary (yes/no) or open-ended. The history $h_{t}=\{q_{1},a_{1},\dots,q_{t-1},a_{t-1}\}$ represents all previous interactions. At each step $i$, the current possibility set $\Omega_{i}$ is updated based on prior interactions. The updated set $\Omega_{i}$ contains all elements consistent with the history $h_{t}$. Questions in the first $\delta*T$ turns are selected strategically by traversing the decision tree of potential questions, and the remaining turns are reserved for making informed decisions about what the outcome is, based on $h_{i}$ and $\Omega_{i}$. Here $\delta$ denotes the proportion of turns reserved for asking information-seeking questions. The process ends when the Questioner identifies $\omega$ or reaches the maximum number of turns $T$.
Upon reaching terminal states in the decision tree (where $|\Omega_{t}|$ $\leq 2$), the Questioner LLM transitions from asking information-seeking questions to making targeted questions about specific outcomes. In the Closed Set scenario, where the target is one of the items within a predefined set $\Omega$, the tree construction is well-defined. However, this space is often unknown in real-world situations, leading to an Open Set scenario, where the models operate without prior knowledge of the outcomes. To address this, we follow the approach used in Hu et al. (2024), where we directly prompt the Question Generator to first define the initial possibility space $\Omega$ based on the problem description and then update it progressively based on the history of interactions $h_{t}$.
In practice, it is important to declare the set of all possible items $\Omega$ in the initial prompt for the Questioner LLM to avoid longer dialogues and failures due to wild guesses in the Closed Set scenario. Hence, to maintain logical consistency, we ensure that the Questioner LLM is explicitly informed about $\Omega$ once at the beginning of each interaction. This helps in constraining the items guessed in the targeting questions. To justify this modification, we report results with and without the $\Omega\text{-}aware$ prompt.
3.2 Decision Tree of Questions
The question generation process employs a hierarchical approach that combines LLMs with Monte Carlo Tree Search (MCTS) to efficiently construct a tree of potential questions represented by nodes. In the information-seeking phase, at each turn $t$, the system either traverses or expands the tree to find the most optimal questions to ask, with the goal of maximizing information gain. This tree is cached and re-used across all data-points belonging to a given dataset.
The LLM-based question generator takes two inputs: the current possibility set $\Omega_{t}$ and the ancestral context $\mathcal{A}_{\Omega t}$ which consists of the sequence of question-answer pairs corresponding to all the ancestor nodes of the current node that led to the formation of $\Omega_{t}$. This context prevents the generation of redundant or previously asked questions, thereby providing a clearer path of deduction. Let $v$ represent a node in the decision tree at turn $t$. With slight abuse of notation, we define $\Omega_{v}$ as the possibility set corresponding to node $v$. For each non-terminal node $v$, where $|\Omega_{v}|>2$, the LLM generates $m$ candidate questions which form its immediate child nodes:
$Q_{v}=\{q_{v}^{(1)},q_{v}^{(2)},...,q_{v}^{(m)}\}=LLM_{gen}(\Omega_{v},\mathcal{A}_{\Omega v})$ (1)
For each generated question, $LLM_{gen}$ performs a binary partitioning of the possibility set $\Omega_{v}$ into two disjoint subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$, corresponding to affirmative and negative responses respectively, such that $\Omega_{v}=\Omega_{v}^{A}\cup\Omega_{v}^{N}$. The LLM is prompted to generate questions that maximize information gain by creating balanced partitions where:
$|\Omega_{v}^{A}|-|\Omega_{v}^{N}|\rightarrow 0$ (2)
Each question spawns two child nodes $v^{A}$ and $v^{N}$ corresponding to affirmative and negative responses, storing $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ respectively. For example, if the possibility set is $\Omega_{v}=\{\text{flu},\text{pneumonia},\text{enteritis},\text{asthma}\}$, the question “Do you have difficulty breathing?” might partition the set as: $\Omega_{v}^{A}=\{\text{pneumonia},\text{asthma}\},\quad\Omega_{v}^{N}=\{\text{flu},\text{enteritis}\}$.
3.3 Uncertainty-Based Reward
To quantify the effectiveness of each question in reducing uncertainty, entropy-based metrics are employed Hu et al. (2024). Following the formulation of reward structure in UoT, let $p_{v}^{A}=|\Omega_{v}^{A}|/|\Omega_{v}|$ and $p_{v}^{N}=1-p_{v}^{A}$. The expected information gain at $v$ is:
$IG_{v}(X)=-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}$ (3)
The reward function, given by $R_{IG}(v)$, achieves its maximum value when the subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ have equal probabilities, signifying the greatest reduction in uncertainty. It attains its minimum value when one of the subsets has a probability of zero, indicating no reduction in uncertainty. To normalize and sharpen rewards, a scaling parameter $\lambda>0$ is employed. Formally, $R_{IG}(v)$ is defined as:
$R_{IG}(v)=\frac{-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}}{1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|}.$ (4)
The expected reward of asking a question, $R_{e}(v)$, is obtained by recursively adding the immediate reward $R_{IG}(v)$ (i.e., information gain at the current node) and the expected rewards of its child nodes. The total information gained until node $v$ of an interaction trajectory is denoted by the accumulated reward, $R_{a}(v)$, which is obtained by starting at the root and propagating down to the node $v$.
[…]
$R_{a}(v):=R_{IG}(v)+\begin{cases}0&v\text{~{}is~{}root},\\ R_{a}(\textit{parent}(v))&\text{otherwise.}\end{cases}$ (6)
where $v^{A}$ and $v^{N}$ are the child nodes of $v$ corresponding to affirmative and negative responses respectively. For each non-terminal response-specific node $v^{\alpha}$, $\alpha\in\{A,N\}$, the expected reward is computed as the average of the expected rewards of its child nodes, denoted by $\textit{children}(v^{\alpha})$. Formally:
$R_{e}(v^{\alpha})=\frac{1}{|\textit{children}(v^{\alpha})|}\sum_{v^{\prime}\in\textit{ children}(v^{\alpha})}R_{e}(v^{\prime})$ (7)
3.4 Monte Carlo Tree Search
The decision tree of questions is constructed iteratively when MCTS is executed over $K$ iterations at each decision-making step across samples. Each iteration includes four phases:
1. Selection: Considering the current node as the root, a child node is selected based on the widely used UCT formulation that balances exploration and exploitation:
$UCT(v)=\frac{R_{\text{total}}(v)}{N_{v}}+C\sqrt{\frac{\ln N_{p}}{N_{v}}}$ (8)
where $R_{a}(v)$ is the accumulated reward, $N_{v}$ is the visit count of node $v$, $N_{p}$ is its parent node’s visit count, and $C$ is the exploration constant.
2. Expansion: The expansion step aims to create child nodes by generating potential questions that split the possibility set, thereby narrowing down the search space of the tree. When the selected node is non-terminal ($|\Omega_{v}|>2$) and does not have child nodes, candidate questions are formulated according to Equation 1. If child nodes already exist, we skip this step and proceed to simulation.
3. Simulation: A rollout policy is used to estimate the expected reward of the selected node by simulating a random interaction up to a predefined depth $d_{s}$ or a terminal state, whichever occurs first. In the rollout phase, we traverse $d_{s}$ levels down from the selected node, choosing one of its child nodes at random, following a single level of expansion if child nodes did not exist. This step provides an estimate of the long-term utility of selecting a particular question.
4. Backpropagation: After reaching a leaf node $u$ at the end of the simulation, the expected reward $R_{e}(u)$ is backpropagated up the tree to update the ancestor nodes along the path to the selected (simulated) node. During backpropagation, the cumulative reward, $R_{\text{total}}(u)$ of every ancestor node of $u$ is updated, and their visit counts are incremented by 1. Formally, for each node $v$ in the path from $u$ to the root:
$R_{\text{total}}(v)\leftarrow R_{\text{total}}(v)+R_{e}(u)$ (9)
This backpropagation process ensures that nodes closer to successful outcomes, i.e., nodes leading to higher rewards, are more likely to be selected in subsequent iterations.
Finally, after completing $K$ iterations of MCTS, the system asks the question having the highest expected information gain. This corresponds to finding $v^{*}$, the child node of the current root having the maximum expected reward:
$v^{*}=\arg\max_{v^{\prime}\in\mathbb{C}(\text{root})}R_{e}(v^{\prime})$ (10)
3.5 Feedback Mechanism and Cluster-Based Reward Adjustment
Learning from experience is a critical component of decision-making systems. We introduce a clustering-based approach for dynamic reward adjustments by propagating feedback through the tree after a successful conversation. When the system successfully identifies the target, it triggers a series of updates to the bonus rewards of nodes that led to the target.
3.5.1 Cluster Assignment and Creation
When a new sample arrives, its text embedding is created and compared with existing cluster medoids to determine its assignment. A similarity threshold $\tau$ is used to decide whether to assign the embedding to an existing cluster or create a new one. Let $\mathbf{e}$ represent the embedding of the current data point and $\mathbf{m}_{k}$ denote the medoid of cluster $k$. The similarity condition is given by:
$\text{Similarity}(\mathbf{e},\mathbf{m}_{k})=\frac{\mathbf{e}\cdot\mathbf{m}_{k}}{\|\mathbf{e}\|_{2}\|\mathbf{m}_{k}\|_{2}}\geq\tau$ (11)
where $\cdot$ represents the dot product, and $\|\cdot\|_{2}$ represents the L2 norm. If no existing cluster satisfies this condition, a new cluster is created with $\mathbf{e}$ as its initial medoid. If assigned to an existing cluster, $\mathbf{e}$ is added to that cluster’s embeddings, and the medoid $\mathbf{m}_{k}$ is recomputed as:
$\mathbf{m}_{k}=\underset{\mathbf{x}\in\mathcal{C}_{k}}{\text{argmax}}\sum_{\mathbf{y}\in\mathcal{C}_{k}}\frac{\mathbf{x}\cdot\mathbf{y}}{\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}}$ (12)
where $\mathcal{C}_{k}$ is the set of embeddings in cluster $k$.
3.5.2 Bonus Rewards for Feedback Propagation
For each node $v$, we maintain a dictionary $B_{k}(v)$ which maps each cluster $k$ to the bonus reward (initialized as zero) earned on reaching the target successfully. Once a target is correctly identified, $B_{k}(v)$ is updated for each node along the path from the current node back to the root. $B_{k}(v)$ is a proportion of the node’s cumulative reward and depends on the absolute depth of the node, given by $d_{v}$ controlled by an exponential decay as we move upward in the tree. The bonus reward for node $v$ corresponding to cluster $k$ is updated after every success:
$B_{k}(v)\leftarrow B_{k}(v)+\beta\cdot R_{\text{total}}(v)\cdot\gamma^{d_{v}}$ (13)
where $\beta$ is a task-specific scaling factor (denoting proportion of total reward), and $\gamma\in(0,1)$ is a decay factor that controls the influence of the bonus, keeping it higher for early-stage questions which have proven effective to narrow down the large possibility set, and lower for questions near the terminal nodes which are often more specific to each case.
To incorporate these bonus rewards into decision-making, we modify the UCT formula used in the selection step to handle similar cases while maintaining exploration capabilities. We introduce a cluster-specific bonus term to bias selection toward nodes associated with successful outcomes for similar data points. The modified UCT formula becomes:
$UCT_{\text{feedback}}(v,k)=\frac{R_{\text{total}}(v)}{N_{v}}+C\sqrt{\frac{\ln N_{p}}{N_{v}}}+B_{k}(v)$ (14)
This modification prioritizes nodes from clusters associated with successful outcomes when processing samples similar to their medoids. Algorithm 1 shows the end-to-end summary of the proposed framework.
Algorithm 1 MISQ-HF
1:Dataset $\mathcal{S}$, question ratio $\delta$, Embedding model, cluster-embeddings hashmap $\mathcal{C}$, similarity threshold $\tau$, maximum turns $T$, MCTS iterations $K$, LLM
2:Initialize $B_{k}(v)=0$, $\forall k\in\{1,2,\text{...},|\mathcal{C}|\}$ for each node $v$
3:for sample $s\in\mathcal{S}$ do
4:     $\mathbf{e}_{s}\leftarrow\text{Embedding}[\text{{description}}(s)]$
5:     $\mathcal{C}_{k}\leftarrow\text{AssignCluster}(\mathbf{e}_{s},\mathcal{C},\tau)$triangleright Use Eq. 11-12
6:     Initialize $\Omega_{v}=\Omega$ at current root node $v$
7:     Initialize conversation history $h=\emptyset$ and $t=0$
8:     while $t<T$ and $\textit{target}(s)\textit{ undetected}$ do
9:         if $t<\delta\cdot T$ and $|\Omega_{v}\geq 2|$ then
10:              $v^{*}\leftarrow\text{MCTS}(v,K,\mathcal{C}_{k})$triangleright Use Eq. 4-9, 14
11:              $q_{t}\leftarrow\textit{question}(v^{*})$triangleright Use Eq. 10
12:              $a_{t}\leftarrow\text{UserResponse}(q_{t},h)$
13:              $v\leftarrow\textit{child}(v^{*}|a_{t})$
14:         else
15:              $q_{t}\leftarrow\text{TargetingPrompt}(\Omega_{v})$
16:              $a_{t}\leftarrow\text{UserResponse}(q_{t},h)$
17:         end if
18:         $h\leftarrow h\cup\{q_{t},a_{t}\}$
19:         $t\leftarrow t+1$
20:     end while
21:     if $\textit{target}(s)\textit{ detected}$ then $\triangleright$ Successful conversation
22:         $v^{\prime}\leftarrow v$triangleright Feedback Propagation
23:         while $\textit{parent}(v^{\prime})\neq\emptyset$ do
24:              Update $B_{k}(v^{\prime})$triangleright Use Eq. 13
25:              $v^{\prime}\leftarrow\textit{parent}(v^{\prime})$
26:         end while
27:     end if
28:end for
Model Method $\Omega$-aware DX MedDG FloDial
[…]
Llama 3.3 70B Instruct UoT $\times$ 72.11 1.54 0.36 79.51 2.09 4.95 34.64 6.84 43.76
MISQ $\times$ 75.00 2.17 0.05 86.56 3.39 0.40 35.29 9.09 3.99
MISQ-HF $\times$ 80.76 1.94 0.21 86.78 3.29 0.78 39.86 9.09 4.07
DP 88.46 3.15 - 84.14 3.93 - 21.56 13.72 -
UoT 79.80 1.65 0.77 89.86 2.16 4.84 60.78 8.47 44.61
MISQ 92.30 1.28 0.48 92.29 3.44 3.59 62.74 9.73 5.16
MISQ-HF 98.07 1.84 0.04 93.39 3.35 0.54 67.97 9.81 3.97
Mixtral 8*7B Instruct DP 50.00 3.50 - 76.43 3.91 - 16.99 14.23 -
UoT 76.92 1.43 0.45 83.70 2.19 5.70 39.21 7.01 45.11
MISQ 63.46 2.63 0.08 76.55 3.33 0.17 47.71 10.45 1.66
MISQ-HF 76.92 2.40 0.06 84.58 3.08 0.33 49.01 9.62 1.46
GPT-4o DP 73.07 3.48 - 81.27 3.98 - 43.79 14.86 -
UoT 82.69 1.18 0.17 88.79 2.03 1.81 59.47 8.14 41.86
MISQ 87.50 1.97 0.03 89.20 3.46 0.60 74.50 10.15 4.10
MISQ-HF 99.03 2.19 0.05 90.30 3.42 0.41 72.54 10.36 2.94
Table 1: Results on Medical Diagnosis and Troubleshooting Domain in a Closed Set scenario. $\Omega$-aware denotes whether the Questioner Model was informed about the possibility space $\Omega$ once at the beginning of the conversation.
4 Experiments
In this section, we provide an overview of our experiments, including descriptions of the datasets, baselines, evaluation metrics, and implementation details.
4.1 Datasets
We aim to demonstrate the effectiveness and versatility of our approach by focusing on three diverse domains: Medical Diagnosis, Troubleshooting, and 20 Questions. We use the following datasets pre-processed by Hu et al. (2024).
In Medical Diagnosis, a patient initially reports a brief description of their symptoms, based on which the doctor asks questions to diagnose the disease. The maximum number of turns $T$ was limited to 6 in the experiments. Two datasets were used. The DX dataset by Xu et al. (2019) contains 104 doctor-patient dialogues and five diseases in its test set. The MedDG dataset, which originally included over 17,000 conversations across 15 disease types, was refined by removing inconsistent samples. We used 454 high-quality samples for evaluation. Open-ended responses are allowed in MedDG to test the system’s generalization capabilities in less constrained scenarios. Both datasets limit interactions to 6 turns. In the Troubleshooting domain, customer support technicians interact with users to identify faults in systems such as cars or electronic devices. We use the FloDial dataset by Raghu et al. (2021), containing 153 dialogues across 153 unique fault types. Maximum number of turns $T$ was limited to 20. In the 20 Questions domain, the task involves identifying a target item by asking up to 20 yes-or-no questions. The Common dataset includes 111 items spanning categories such as animals, places, food, and objects, and the Things dataset Hebart et al. (2019), was filtered to include 200 items.
4.2 Baselines
The proposed system is compared with two primary baselines: Direct Prompting (DP) and the Uncertainty of Thoughts (UoT) framework Hu et al. (2024). DP involves querying the LLM directly to generate the next question without leveraging any structured planning mechanism. This approach evaluates the basic capability of the LLM to ask effective questions without additional search strategies. It serves as a minimal-effort benchmark, highlighting the improvements achieved by strategic decision-making approaches. UoT implements a tree-based planning method that performs an exhaustive expansion of all possible question paths to maximize information gain. While UoT demonstrates strong performance in reducing uncertainty and improving task success rates, its reliance on full tree exploration up to a predefined depth can lead to significant computational overhead, particularly within large possibility spaces. This baseline is included to measure the improvement in efficiency achieved by the proposed MISQ-HF approach, which selectively expands promising branches rather than exploring all possibilities exhaustively. Comparisons with methods such as Chain-of-Thought (CoT) Wei et al. (2022) and Tree-of-Thoughts (ToT) Yao et al. (2023) are omitted, as prior work by Hu et al. (2024) has demonstrated that UoT consistently outperforms these techniques in both efficiency and effectiveness.
4.3 Evaluation
We use three key metrics to assess both the effectiveness and efficiency of the question-asking system. First, we measure Success Rate (SR), defined as the percentage of cases where the system correctly identifies the target item within the maximum allowed turns. Second, we track Mean Conversation Length in Successful Cases (MSC), which represents the average number of turns needed to reach the correct answer (target) in successful interactions. Finally, we introduce a novel metric to track Question Generation Calls (QGC), defined as the number of LLM prompting calls required for question generation. Lower values indicate better computational efficiency. This metric directly measures the system’s ability to minimize expensive LLM operations while maintaining performance. Together, these metrics provide a comprehensive view of the system’s performance, balancing success rate, conversation efficiency, and computational overhead.
4.4 Experimental Setup
Models. We employed three different LLMs of varying sizes as the Questioner. Llama 3.3 70B Instruct Grattafiori et al. (2024) and Mixtral 8*7B Instruct Jiang et al. (2024) were accessed via the AWS Bedrock API (2023). GPT-4o was accessed via API provided by OpenAI et al. (2024). The user (Answerer) was simulated by Llama 3.3 70B Instruct in all tasks. We prompted the model with ground truth details (e.g., the patient’s disease or fault description) and maintained a separate conversation history. The temperature was set to 0.
MCTS implementation. We set the number of iterations $K=10$ and exploration constant $C=0.2$. The maximum simulation depth $d_{s}$ was set to 3 to balance computational efficiency with search effectiveness. For each $\Omega_{v}$, the question generator creates $m=3$ potential questions to maintain diversity. For the reward calculation in $R_{IG}(v)$, the scaling parameter $\lambda$ was set to 0.4, for both MCTS and UoT.
Feedback mechanism. We used a decay factor $\gamma=0.9$ for the bonus rewards, ensuring early-stage questions receive higher weights. The cluster similarity threshold $\tau$ was set to 0.9 in terms of cosine similarity, and the bonus scaling factor $\beta$ was set to 0.2 for all tasks. Problem descriptions were embedded using DistilBERT Sanh et al. (2019) for the troubleshooting domain, and Clinical-BERT Wang et al. (2023) for medical diagnosis. To optimize computational efficiency, we cache the decision tree, allowing reuse across all samples within each dataset. The system tracks Mean QGC through a thread-safe counter.
5 Results and Discussion
Table 1 shows our results in the Closed Set scenario. Our approach requires fewer QGC while maintaining higher SR compared to baseline methods, indicating both improved effectiveness and computational efficiency. Notably, the $\Omega$-aware Questioner consistently performs better as compared to when unaware. To demonstrate the importance of the feedback mechanism, we also report results using MISQ, our framework without the hierarchical feedback component, to rationalize the design choice and support an ablation study.
In the Medical Diagnosis domain, MISQ-HF achieves superior performance across both datasets, with an average reduction of 6.77 times in Mean QGC across all three LLMs. Using Llama 3.3 70B as a Questioner, MISQ-HF required only 0.04 Mean QGC compared to UoT’s 0.77 on DX, showing a 19.25 times reduction. Similarly, a reduction of 8.97 times is observed on MedDG. This efficiency stems from selective tree expansion rather than exhaustive exploration. In the medical domain, the cost of misdiagnosis far outweighs the overhead of a few additional clarifying questions. While MSC is marginally higher for MISQ-HF compared to UoT, this trade-off can be overseen by the higher success rate of diagnosis, which directly impacts patient outcomes.
In the Troubleshooting domain using the FloDial dataset, MISQ-HF continues to excel with higher SR, surpassing UoT by 20.68% improvement on average across the three LLMs. Our framework also achieves notable computational efficiency as compared to UoT, with an average of 18.63 times reduction in Mean QGC. While troubleshooting conversations typically require more turns due to fault complexity, MISQ-HF maintains a similar MSC (a difference of less than 3 turns on average) while delivering improved performance. With GPT-4o, MISQ-HF achieved an SR of 72.54% with 27 clusters, the largest consisting of 31 samples.
Table 2 shows performance on the 20 Questions domain, evaluated using the Common and Things datasets. Due to a lack of problem description or initial hints about the target, we do not evaluate MISQ-HF on this domain. For both datasets, MISQ consistently outperforms UoT and DP in all three metrics. These results demonstrate the scalability of the proposed approach to larger possibility spaces.
Method $\Omega$- aware Common Thing
[…]
Llama 3.3 70B Instruct
UoT $\times$ 39.63 8.27 4.08 19.00 9.78 4.48
MISQ $\times$ 41.44 8.43 5.05 23.5 9.57 1.57
DP 45.94 13.70 - 32.50 13.27 -
UoT 61.26 9.94 7.92 35.50 11.43 3.40
MISQ 74.77 9.90 4.74 59.50 10.68 3.31
Mixtral 8*7B Instruct
DP 8.10 14.33 - 7.50 13.46 -
UoT 28.82 11.56 4.34 12.50 13.52 5.91
MISQ 37.83 11.38 2.39 20.00 11.50 0.06
GPT-4o
DP 63.06 14.72 - 40.50 14.16 -
UoT 74.77 8.59 5.88 47.00 9.13 2.75
MISQ 85.58 8.51 4.86 55.50 9.54 2.19
Table 2: Results on 20 Questions Datasets in the Closed Set scenario
The results in the Open Set scenario are demonstrated on two datasets where the range of possible outcomes is often unpredictable and varies in size. Problem descriptions were used to initialize $\Omega$ to a set of 5 possibilities, and it was progressively updated based on conversation history. We used FloDial for troubleshooting, which has a larger possibility space, and DX for medical diagnosis, which has a smaller possibility space. Table 3 shows consistently lower Mean QGC while maintaining similar SR in both domains.
Method FloDial DX
[…]
DP 16.99 14.80 - 29.80 3.22 -
UoT 28.10 7.56 11.42 35.57 2.35 10.22
MISQ 28.10 8.02 5.33 36.53 2.73 6.52
MISQ-HF 28.75 6.95 5.10 37.50 2.35 6.32
Table 3: Results in the Open Set scenario. Possibility set $\Omega$ is unknown to the Questioner LLM.
[…]
The frequency of QGC for exhaustive tree search and MCTS differs significantly. An LLM call creates $m$ questions, each splitting $\Omega$ into affirmative and negative response sets, thereby giving rise to $2m$ nodes. For UoT, the first simulation requires $1+2m+(2m)^{2}+\dots+(2m)^{d_{s}}=\frac{(2m)^{d_{s}+1}-1}{2m-1}$ calls, and $(2m)^{d_{s}}$ calls for subsequent turns if paths are expanded till $d_{s}-1$, where $d_{s}$ is the depth-limit for simulation. For MCTS, the first expansion requires 1 call, and simulation for selecting a question at each turn requires $K\cdot d_{avg}$, where $K$ is the number of iterations and $0\leq d_{avg}\leq d_{s}$ because traversing requires 0 LLM calls when the nodes are already expanded. Exhaustive search scales exponentially with $d_{s}$, while MCTS scales linearly, highlighting its efficiency for deeper searches in the simulation phase.
Limitations. The system currently does not incorporate mechanisms to learn from mistakes in failure cases, which could potentially refine its decision-making process. Explicitly penalizing suboptimal or redundant questions also requires careful design of reward function and remains an open challenge. The framework can benefit from incorporating confidence metrics to better quantify uncertainty in risk-sensitive decision making domains. Addressing these limitations is a promising direction for future work.
6 Conclusion
In this paper, we introduced Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback (MISQ-HF), a novel framework for adaptive question-asking with a goal to optimize conversational efficiency and accuracy. By leveraging selective tree expansion and cluster-based feedback, our approach reduces computational overhead while maintaining high success rates across diverse domains such as medical diagnosis, troubleshooting, and general information-seeking tasks. We demonstrate the potential of combining structured and efficient planning with historical learning to dynamically adapt question-asking strategies. Our contributions aim to refine the decision-making process, enabling conversational AI systems to handle increasingly complex and open-ended scenarios with better planning.
Acknowledgments
This work is supported by the National Science Foundation (NSF) award IIS-2336769. The authors thank Hafsah Usman, student of the University of Washington, Seattle, for her valuable assistance in running a part of the experiments.
Appendix A Conversation Prompts for 20 Questions
A.1 Questioner Prologue
Let us play the game of 20 questions. I am impersonating the thing, X. {inform_set} You will ask me up to 20 questions which start with ‘Is X’ and can only be answered by yes or no, and I will answer each one truthfully based on being X. Let us begin. Ask me the first question.
{inform_set} “X is possibly one of the following: $\Omega$
This is given only once at the beginning of the conversation.
A.2 Answerer/User Simulator Prologue
Let us play the game of 20 questions. You are the answerer and I am questioner. X is ‘{target_item}’. I will ask you up to 20 questions and you should answer each one truthfully based on being X, by saying Yes or No.
Note that you must never reveal X, until I guess it correctly.
If I guess X correctly in my question, directly respond “You guessed it. X is ‘{target_item}’.” instead of saying yes. Let us begin. Here is my first question.
{target_item} the ground truth target item based on which user responds
A.3 Targeting Prompt
Note that you should guess and ask what X exactly is from now on. X is possibly one of the following: {item_set_terminal}, or other from the initial set of possibilities. Use the past conversation for context to avoid redundant questions. The question must start with ‘Is X …’
{item_set_terminal} $\Omega_{v}$ (the set of items remaining at leaf node $v$)
Target Question: Is X ‘{item}’?
A.4 Question Generation Prompt
Here are all the X: {item_set}
Design a question about X that can only be answered by Yes or No. Then classify the possible X above based on this question. If the answer is ’YES’, put this X into ’YES: …’, otherwise to ’NO: …’. Finally calculate how many X in YES and NO. Notably, this question should fulfill that the count of YES and NO are almost the same with a permissible discrepancy of no more than one!
{ancestral_context}
Based on this information, create most relevant {m} questions to classify the above X correctly. Your response should strictly follow the template:
Question 1: Is X …?
YES: comma-separated, list of things, …
Count of YES: …
NO: comma-separated, list of things, …
Count of NO: …
{item_set} $\Omega_{v}$ (the set of items remaining at node $v$)
{ancestral_context} “For context, following questions were already asked to build the above set of possibilities: Question? No; Question? Yes; Question? Yes…” if ancestral question nodes exist, else blank string.
{m} number of potential questions to generate (actions)
Appendix B Conversation Prompts for Medical Diagnosis
B.1 Questioner Prologue
You are a doctor, and your patient self-reports that: {problem_description}. {inform_set} You should ask your patient questions in English about symptoms which can only be answered by Yes or No, in order to find what disease this patient suffers from. Use the ongoing conversation for context to avoid redundant questions. Let us begin. Ask me the first question.
B.2 Answerer/User Prologue
You are the patient suffering from ‘{target_item}’, and I am the doctor. I will ask you up to 6 questions, and you should answer each one truthfully based on your disease, by saying Yes or No. Note that you must never reveal the disease until I tell it correctly. If I tell the disease correctly in my question, directly respond: “You are right. I am experiencing ‘{target_item}’.” Let us begin. Here is my first question.
B.3 Question Generation Prompt
You are a doctor. Here are all the possible diseases that the patient may suffer from: {item_set}
Design a question to ask your patient regarding symptoms of their illness that can only be answered by Yes or No. Then classify the possible diseases above based on this question. If the answer is ’YES’, put this disease into ’YES: …’, otherwise to ’NO: …’. Finally calculate how many diseases are in YES and NO. Notably, this question should fulfill that the count of YES and NO are almost the same with a permissible discrepancy of no more than one!
{ancestral_context}
Based on this information, create most relevant {n} questions to ask (and classify the above diseases). Your response should strictly follow the template:
Question 1: …?
YES: comma-separated, list of disease names, …
Count of YES: …
NO: comma-separated, list of disease names, …
Count of NO: …
B.4 Targeting Prompt
Note that you should point out and ask what disease the patient suffers from now.
The patient is probably suffering from one of the diseases below: {item_set_terminal}, or other from the initial set of possibilities.
Refer to the past conversation regarding the patient’s symptoms. Never repeat previously asked questions. Respond with 1 new question only, following the format: “Are you experiencing [disease name]?” Ensure that [disease name] was not asked before.
Target Question: Are you experiencing ‘{item}’?
B.5 Classify Prompt
Here are all diseases that the patient may suffer from: {item_set}
{problem_description} For each disease under this report, if the patient is possibly affected, put this disease into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
YES: comma-separated, list of disease names, …
NO: comma-separated, list of disease names, …
B.6 Open Set Prompts
Initial Prompt:
You are a doctor, and your patient self-reports that: {problem_description}. Please propose {size} diseases that you think your patient may suffer from. Your response should only be a list like: [”disease1”, ”disease2”, …]
Renewal Prompt:
Based on the conversation history, please propose {size} diseases that your patient may suffer from. The list of {size} diseases should contain {existing_items}. Your response should only be a list like: [”disease1”, ”disease2”, …]
Appendix C Conversation Prompts for Troubleshooting
C.1 Questioner Prologue
You are a technician, and your client self-reports that: {problem_description}. {inform_set} You should ask your client questions with specific situations which can only be answered by Yes or No, in order to find which issue this client is facing. Use the ongoing conversation for context to avoid redundant questions. Let us begin. Ask me the first question.
C.2 Answerer/User Prologue
You are the client with a device that has ‘{target_item}’, and I am the technician. I will ask you up to 20 questions, and you should answer each one truthfully based on the issue of your device, by saying Yes or No. Note that you must never reveal the issue name until I tell it correctly. If I tell your issue correctly in my question, directly respond: “You are right. My device has issues with ‘{target_item}’.” Let us begin. Here is my first question.
C.3 Question Generation Prompt
You are a technician. Here are all the issues that the client may face: {items_set}
Design a question to ask your client with a specific situation that can only be answered by YES or NO. Then classify the possible issues above based on this question. If the answer is ’YES’, put this issue into ’YES: …’, otherwise to ’NO: …’. Finally calculate how many issues are in YES and NO. Notably, this question should fulfill that the count of YES and NO are almost the same with a permissible discrepancy of no more than one!
{ancestral_context}
Based on this information, create the most relevant {n} questions to classify the above issues correctly. Your response should strictly follow the template:
Question 1: …?
YES: comma-separated, list of issue names, …
Count of YES: …
NO: comma-separated, list of issue names, …
Count of NO: …
C.4 Targeting Prompt
Note that you should now point out and ask what issue the client is facing. Refer to the past conversation for context and never ask redundant questions. Respond with 1 new question only, following the format: “Are you experiencing [issue name]?” Ensure that [issue name] was not asked before.
Target Question: Are you experiencing ‘{item}’?
C.5 Classify Prompt
Here are all the issues that the client may face: {item_set}
{problem_description} For each issue under this report, if the client is possibly facing it, put this issue into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
YES: comma-separated, list of issue names, …
NO: comma-separated, list of issue names, …
C.6 Open Set Prompts
Initial Prompt:
You are a technician, and your client self-reports that: {problem_description}. Please propose {size} issues that you think your client may face with. Your response should only be a list like: [”issue1”, ”issue2”, …]
Renewal Prompt:
Based on the conversation history, please propose {size} issues that your client may face with. The list of {size} issues should contain {existing_items}. Your response should only be a list like: [”issue1”, ”issue2”, …]
Appendix D Examples
Two sequences of history are maintained respectively for Questioner LLM and User (Answerer LLM). Each stores the questions asked and answers given so far. At the beginning of the conversation, the Question LLM is given the Questioner Prologue as the first prompt, and the Answerer LLM is given the Answerer Prologue as the first prompt. The rest of the conversation history remains the same, except for the switching between the “system” role and the “user” role for each LLM. Table 4, 5, 6 show example conversations from three domains.
S.No. System User
Problem Description:
Hello doctor, I have been in this situation for several years. Every summer, I often have diarrhea, but it has gotten worse in the past year or two. There is no regular pattern, sometimes it happens every two to three days when it’s mild, and two to three times a day when it’s severe. $\dots$ I went to a major hospital for examination before, and they found no parasite eggs or fungi, just a slight imbalance in the bacterial colonies, but the medication didn’t work. I don’t know what my condition is, I wonder if any doctor can provide some answers. Thank you. (Male, 28 years old).
1 Do you experience any symptoms related to the respiratory system, such as coughing or difficulty breathing? No, I don’t experience any symptoms related to the respiratory system, such as coughing or difficulty breathing. My main issue is with my stomach and bowel movements. I have been experiencing diarrhea, stomach aches, and urgent bowel movements, especially during the summer. The stool is often loose, sticky, and not well-formed, and it’s difficult to flush. I’ve been having this issue for several years, and it’s been getting worse over the past year or two.
2 Do you experience any abdominal pain or discomfort? Yes, I do experience abdominal pain and discomfort. It usually starts with my stomach aching more and more, and then I have to go to the bathroom. The pain can be quite severe, especially when the symptoms are bad. Sometimes, the stool is not very loose, but it’s accompanied by mucus and is not well-formed, which makes it difficult to flush. The abdominal pain and discomfort are quite frequent, and it’s been a persistent issue for me.
3 Do you experience any symptoms related to digestion, such as nausea or vomiting? I don’t usually experience nausea or vomiting, but I do have diarrhea, which can be quite loose and sticky at times. The stool is often not well-formed and can be difficult to flush. The frequency of the diarrhea can vary, sometimes it’s every two to three days when it’s mild, and sometimes it’s two to three times a day when it’s severe. The symptoms can be quite unpredictable and irregular, which makes it challenging to manage.
4 Are you experiencing symptoms of irritable bowel syndrome? You are right. I have Irritable bowel syndrome.
Table 4: An example conversation from MedDG dataset using MISQ-HF with GPT-4o.
S.No. System User
Problem Description:
My car continues to overheat. The hood is so hot that I think I could melt cheese on the hood! Please help.
1 Is the issue related to the vehicle’s braking system? No.
2 Is the issue related to a power or electrical problem? No.
3 Is the issue related to overheating or temperature control? Yes.
4 Is the issue related to a mechanical part? Yes.
5 Is there any visible fluid leaking from the vehicle? Yes.
6 Are you experiencing a radiator leak? You are right. It’s a Radiator leak.
Table 5: An example conversation from FloDial dataset using MISQ-HF with GPT-4o.
S.No. System User
1 Is X a living organism? Yes.
2 Is X an animal that primarily lives in water? No.
3 Is X a type of food? No.
4 Is X a mammal? No.
5 Is X capable of flight? No.
6 Is X a reptile? Yes.
7 Is X capable of changing its skin color for camouflage? Yes.
8 Is X a ’Chameleon’? You guessed it. X is Chameleon.
Table 6: An example conversation from Common dataset using MISQ with GPT-4o.

### v2

@@
 Abstract
-The ability to identify and acquire missing information is a critical component of effective decision making and problem solving.
-With the rise of conversational artificial intelligence (AI) systems, strategically formulating information-seeking questions becomes crucial and demands efficient methods to guide the search process.
-We introduce a novel approach to adaptive question-asking through a combination of Large Language Models (LLM) for generating questions that maximize information gain, Monte Carlo Tree Search (MCTS) for constructing and leveraging a decision tree across multiple samples, and a hierarchical feedback mechanism to learn from past interactions.
-We present two key innovations: (1) an adaptive MCTS algorithm that balances exploration and exploitation for efficient search over potential questions; and (2) a clustering-based feedback algorithm that leverages prior experience to guide future interactions.
-Each incoming sample is assigned to a cluster based on its semantic similarity with previously observed samples.
-Our UCT (Upper Confidence bound for Trees) formulation selects optimal questions by combining expected rewards, a function of information gain, with a cluster-specific bonus that decays with depth, to emphasize the importance of early-stage questions that have proven effective for narrowing the solution space in similar samples.
-Experiments across three domains, including medical diagnosis and troubleshooting, demonstrate that our method leads to an average of 12% improvement in success rates and a 10x reduction in the average number of LLM calls made per conversation for the search process, in comparison to the state of the art.
-[^1] https://github.com/harshita-chopra/misq-hf
+Effective decision-making and problem-solving in conversational systems require the ability to identify and acquire missing information through targeted questioning.
+A key challenge lies in efficiently narrowing down a large space of possible outcomes by posing questions that minimize uncertainty.
+To address this, we introduce a novel framework that leverages Large Language Models (LLMs) to generate information-seeking questions, with Monte Carlo Tree Search (MCTS) to strategically select questions that maximize information gain, as a part of inference-time planning.
+Our primary contribution includes a hierarchical feedback mechanism that exploits past interaction patterns to guide future strategy.
+Specifically, each new problem is mapped to a cluster based on semantic similarity, and our UCT (Upper Confidence bound for Trees) formulation employs a cluster-specific bonus reward to prioritize successful question trajectories that have proven effective for similar problems in the past.
+Extensive empirical evaluation across medical diagnosis and technical troubleshooting domains shows that our method achieves an average of 12% improvement in success rates and about 10x reduction in the number of LLM calls made for planning per conversation, compared to the state of the art.
+An additional 8% gain in success rate is observed on average when we start with a constrained set of possibilities.
+Our results underscore the efficacy of feedback-aware MCTS in enhancing information-seeking in goal-oriented dialogues.
+[^1] Our code is available here.
 1 Introduction
-The field of conversational AI has been witnessing significant advancements in recent years, with a growing focus on developing systems capable of adaptive and efficient information-seeking strategies.
-As Large Language Models (LLMs) continue to push the boundaries of natural language understanding and generation, researchers are exploring novel approaches to enhance their decision-making capabilities in interactive contexts.
-One of the core challenges in a conversational system that employs multiple rounds of Q&A to arrive at an answer is to determine the next question to ask.
-A common solution to this challenge is to construct a contextual decision tree that allows the system to strategically plan which questions to pose and in what order, refining the process to ensure the correct answer or goal is achieved efficiently.
-LLMs have emerged as powerful tools for planning using trees Zhao et al.
-(2024); Putta et al.
-(2024); Hu et al.
-(2024).
-Notably, the Tree of Thoughts (ToT) Yao et al.
-(2023) approach has demonstrated the potential of leveraging hierarchical structures to improve problem-solving and reasoning in language models.
+Recent advances in conversational AI have transformed human-machine interactions, enabling more sophisticated goal-oriented dialogue systems.
+A fundamental requirement in these systems is the ability to identify and seek out missing information efficiently.
+When confronted with an initially under-specified problem, conversational agents must interact over multiple turns and strategically ask questions that reduce uncertainty while minimizing interaction overhead.
+This necessitates a principled approach to sequential decision-making wherein each question posed must maximize information gain and narrow down the solution search space.
+Large Language Models (LLMs) have shown strong capabilities in natural language understanding and generation, offering a promising basis for enhancing these information-seeking strategies.
+However, effectively applying LLMs to domains that involve complex problem-solving still requires overcoming challenges in efficient planning, exploration-exploitation trade-offs, and adaptation to domain-specific interaction patterns.
+LLMs have emerged as powerful tools for planning using trees [22, 12, 7].
+Notably, the Tree of Thoughts (ToT) [20] approach has demonstrated the potential of leveraging hierarchical structures to improve problem-solving and reasoning in language models.
 However, expanding the full tree can be computationally inefficient and expensive, especially in domains with large search spaces.
-One of the fundamental problems lies in designing systems capable of dynamically adapting their interaction strategy to converge on a specific solution efficiently.
-To address these challenges, Monte Carlo Tree Search (MCTS) has emerged as a promising technique for balancing exploration and exploitation in decision-making processes.
-MCTS has been successfully applied in various domains, from game-playing AI Silver et al.
-(2016, 2018) to robotics Dam et al.
-(2022), and is now being leveraged in natural language processing tasks.
+A fundamental challenge lies in designing systems capable of dynamically adapting their interaction strategy to converge on a specific solution efficiently.
+To address these challenges, Monte Carlo Tree Search (MCTS) [2] has emerged as a promising technique for balancing exploration and exploitation in decision-making processes.
+It has been successfully applied in various domains, from game-playing AI [15, 16] to robotics [3], and is now being leveraged in natural language processing tasks.
 Figure 1: Overview of MISQ-HF.
-The system combines Monte Carlo Tree Search with hierarchical feedback for efficient information seeking.
-Green nodes ( ) and red nodes ( ) represent $v^{A}$ and $v^{N}$ respectively, corresponding to possibilities after affirmative and negative answers to a question.
+In the MCTS block, green nodes ( ) and red nodes ( ) represent $v^{A}$ and $v^{N}$ respectively, corresponding to possibilities after affirmative and negative answers to a question.
 The purple arrows indicate the information-seeking question loop (steps 4-8).
 Towards the end of decision-making, targeted questions are generated, denoted by the loop of dashed arrows (steps 9-6-7-8).
-Finally, when the user response $a_{t}$ (step 7) confirms the target detection, the conversation is terminated (step 10) and the feedback mechanism updates the reward bonus along the successful questioning trajectory in the decision tree, enabling the system to adapt its question selection strategy for similar future cases.
+Finally, when the user response $a_{t}$ (step 7) confirms the target detection, the conversation is terminated (step 10) and the feedback mechanism updates the reward bonus along the successful questioning trajectory in the decision tree, enabling the system to adapt for similar future cases.
 Building upon this foundation, our work introduces a novel framework that combines MCTS with a hierarchical feedback mechanism to achieve adaptive and efficient information-seeking.
-Inspired by prior work on uncertainty-aware planning, our approach extends these ideas with three key contributions:
+Inspired by prior work on uncertainty-aware planning, our approach presents three key contributions:
   1. 1.
-    MCTS on Decision Tree for Question Generation: Traditional tree-based approaches can struggle with the vast search spaces.
-By employing MCTS, our system efficiently constructs and explores promising branches of the decision tree without exhaustive expansion.
-We propose a variant of MCTS that optimizes question selection by considering both immediate and long-term information gain.
-We use the standard Upper Confidence Bound for Trees (UCT) and propose a modified variant that assigns higher weights to questions that have effectively narrowed the solution space in past interactions.
+    MCTS for Question Selection: LLMs can generate multiple valid questions to narrow search spaces, but identifying the optimal ones is crucial.
+Tree-based approaches are known for comprehensive planning.
+By employing MCTS, our system efficiently constructs and explores promising questions in the decision tree without exhaustive expansion, focusing on both immediate and long-term information gain.
+Our modified variant of the Upper Confidence Bound for Trees (UCT) assigns higher weights to questions that have effectively narrowed the solution space in past interactions.
   2. 2.
-    Cluster-Based Feedback Mechanism: We leverage similarity-based clustering based on text embeddings of the initial problem description given by the user.
-We allow the system to learn from historical successes by propagating cluster-specific bonus rewards through the decision tree after a successful interaction.
-This bonus reward decays with depth, recognizing that early-stage questions play a crucial role in narrowing down the solution space by targeting high-level information that can effectively eliminate large subsets of possibilities.
-These questions are often broadly applicable across multiple cases, making them valuable for generalizing strategies.
-In contrast, later-stage questions tend to be more specific to individual cases, focusing on finer details that depend on the unique context of each interaction.
+    Cluster-Based Feedback Mechanism: We cluster similar cases using text embeddings derived from the user’s initial problem description.
+After a successful interaction, the system propagates cluster-specific bonus rewards through the decision tree.
+This bonus reward decays with depth, emphasizing the importance of early-stage questions that effectively reduce the solution space by targeting broad, generalizable information.
   3. 3.
     Efficiency through reduced LLM calls: Our method minimizes computational overhead by strategically limiting LLM interactions during tree construction and simulation.
-We maintain a single decision tree for a given dataset which expands as more data points stream in.
+We maintain a single decision tree for a given dataset, which expands as more data points stream in.
 We also employ lightweight rollout policies limited by predefined depth during the simulation step of MCTS.
@@
 We evaluate our approach across diverse conversational tasks.
-Results demonstrate that our system outperforms existing baselines in both task efficiency and relevance, particularly in scenarios requiring complex reasoning and hierarchical decision-making.
-Additionally, we also highlight the individual contributions of depth-aware MCTS and cluster-based feedback in enhancing the system’s performance.
+Results demonstrate that our system outperforms existing baselines in both task success and efficiency in scenarios requiring complex reasoning and hierarchical decision-making.
+We also highlight the individual contributions of depth-aware MCTS and cluster-based feedback in enhancing the system’s performance.
 2 Related Work
 The integration of LLMs with tree search algorithms has proven effective for enhancing multi-step reasoning and decision-making.
-Guez et al.
-(2018) introduced MCTSnets, combining tree search with neural networks for tasks like Sokoban.
+Guez et al.[5] introduced MCTSnets, combining tree search with neural networks for tasks like Sokoban.
 While it focuses on discrete planning, our approach extends MCTS to hierarchical conversations, optimizing question selection with depth-aware strategies and semantic embeddings.
-Yu et al.
-(2023) used MCTS for dialogue planning with LLM-based simulations, but their approach differs in using open-loop search without feedback mechanisms and focuses only on persuasion tasks.
-Hui and Tu (2024) introduced the Reflection on Search Trees (RoT) framework for generating state-specific guidelines to improve search performance.
-Unlike RoT, which focuses on reasoning efficiency, our approach maintains a single decision tree while incorporating cluster-specific rewards to guide question selection.
-Zhu et al.
-(2021) framed retrieval and answering as a partially observed Markov decision process to adaptively gather evidence, though it does not address learning from feedback in conversational settings.
-In the clinical domain, Li et al.
-(2024) proposed MEDIQ, an adaptive framework leveraging LLMs to identify missing information and ask follow-up questions.
-While MEDIQ emphasizes reliability in high-stakes contexts, our work generalizes by prioritizing early-stage questions to effectively narrow the solution space.
-Xie et al.
-(2024) enhanced LLM reasoning through iterative preference learning inspired by AlphaZero, using MCTS to collect step-level preferences.
-This aligns with our approach of integrating cluster-specific success bonus rewards into MCTS.
-Zhang et al.
-(2024) explored autonomous tree search capabilities in LLMs, demonstrating self-guided reasoning, while our method emphasizes leveraging historical feedback through clustering mechanisms.
-Our work draws inspiration from the Uncertainty of Thoughts (UoT) algorithm introduced by Hu et al.
-(2024), which enhances LLMs’ ability to actively seek information by asking effective questions.
-UoT combines a simulation approach with uncertainty-based rewards motivated by information gain, and a reward propagation scheme to select optimal questions, demonstrating significant improvements in task performance across various domains and baselines like Chain of Thoughts Wei et al.
-(2022) and Tree of Thoughts Yao et al.
-(2023).
-Building upon these principles, our approach extends the concept by integrating a depth-aware MCTS algorithm that optimizes question selection in hierarchical conversations, and a cluster-based feedback system that leverages historical successes to guide future interactions by adding a depth-aware bonus reward to prioritize early-stage questions.
-By integrating these components, our approach achieves a dynamic balance between efficiency and adaptability.
+Yu et al.[21] applied MCTS for dialogue planning with LLM-based simulations, but their method differs by using open-loop search without feedback mechanisms and focusing solely on persuasion tasks.
+Hui et al.[8] proposed the Reflection on Search Trees (RoT) framework for generating state-specific guidelines to improve search performance.
+Unlike RoT, which prioritizes reasoning efficiency, our approach maintains a single decision tree while incorporating cluster-specific rewards to guide question selection.
+Zhu et al.[23] framed retrieval and answering as a partially observed Markov decision process to adaptively gather evidence but does not address learning from feedback in conversational settings.
+In the clinical domain, Li et al.[10] introduced MEDIQ, an adaptive framework leveraging LLMs to identify missing information and ask follow-up questions.
+While MEDIQ focuses on reliability in high-stakes contexts, our method generalizes by prioritizing early-stage questions to narrow the solution space effectively.
+Our work draws inspiration from the Uncertainty of Thoughts (UoT) algorithm [7], which enhances LLMs’ ability to seek information through effective questioning.
+UoT uses uncertainty-based rewards driven by information gain and a reward propagation scheme to optimize question selection, improving performance across various domains and baselines like Chain of Thoughts [18] and Tree of Thoughts [20].
+Extending these principles, we integrate a depth-aware MCTS algorithm for question selection in hierarchical conversations and a cluster-based feedback system that leverages historical successes.
+The depth-aware bonus reward prioritizes early-stage questions, balancing efficiency and adaptability.
 3 Methodology
-In this section, we present our problem formulation, construction of the decision tree, reward function, our MCTS algorithm, and our hierarchical feedback approach.
 3.1 Problem Formulation
 The task is modeled as a sequential interaction between a Questioner (an LLM) and an Answerer (a human; here, simulated by an LLM).
-Let $\Omega$ be the space of possible target items, with an unknown target $\omega\in\Omega$.
+Let $\Omega$ be the space of possible target outcomes, with an unknown target $\omega\in\Omega$.
 The interaction proceeds over turns $t=1,\dots,T$, where $T$ is the maximum allowed turns.
@@
 The history $h_{t}=\{q_{1},a_{1},\dots,q_{t-1},a_{t-1}\}$ represents all previous interactions.
-At each step $i$, the current possibility set $\Omega_{i}$ is updated based on prior interactions.
-The updated set $\Omega_{i}$ contains all elements consistent with the history $h_{t}$.
-Questions in the first $\delta*T$ turns are selected strategically by traversing the decision tree of potential questions, and the remaining turns are reserved for making informed decisions about what the outcome is, based on $h_{i}$ and $\Omega_{i}$.
-Here $\delta$ denotes the proportion of turns reserved for asking information-seeking questions.
-The process ends when the Questioner identifies $\omega$ or reaches the maximum number of turns $T$.
+At each step $t$, the current possibility set $\Omega_{t}$ is updated based on prior interactions.
+The updated set $\Omega_{t}$ contains all elements consistent with the history $h_{t}$.
+Questions in the first $\delta*T$ turns, where $\delta\in(0,1)$, are selected strategically by traversing the decision tree of potential questions, and the remaining turns are reserved for making informed decisions about what the outcome (target) is, based on $h_{t}$ and $\Omega_{t}$.
+Here, $\delta$ denotes the proportion of turns reserved for asking information-seeking questions.
+The process ends when the Questioner identifies $\omega$ or reaches the maximum turns $T$.
 Upon reaching terminal states in the decision tree (where $|\Omega_{t}|$ $\leq 2$), the Questioner LLM transitions from asking information-seeking questions to making targeted questions about specific outcomes.
@@
 However, this space is often unknown in real-world situations, leading to an Open Set scenario, where the models operate without prior knowledge of the outcomes.
-To address this, we follow the approach used in Hu et al.
-(2024), where we directly prompt the Question Generator to first define the initial possibility space $\Omega$ based on the problem description and then update it progressively based on the history of interactions $h_{t}$.
-In practice, it is important to declare the set of all possible items $\Omega$ in the initial prompt for the Questioner LLM to avoid longer dialogues and failures due to wild guesses in the Closed Set scenario.
-Hence, to maintain logical consistency, we ensure that the Questioner LLM is explicitly informed about $\Omega$ once at the beginning of each interaction.
-This helps in constraining the items guessed in the targeting questions.
+To address this, we follow the approach used in [7], where we directly prompt the Question Generator to first define the initial possibility space $\Omega$ based on the problem description and then update it progressively based on the history of interactions $h_{t}$.
+In practice, it is important to declare the set of all possible items $\Omega$ in the initial prompt for the Questioner LLM to avoid longer dialogues and failures due to arbitrary guesses in the Closed Set scenario.
+Hence, we ensure that the Questioner LLM is explicitly informed about $\Omega$ once at the beginning of each interaction.
+This helps fixing the items that appear in the targeting questions.
 To justify this modification, we report results with and without the $\Omega\text{-}aware$ prompt.
 3.2 Decision Tree of Questions
-The question generation process employs a hierarchical approach that combines LLMs with Monte Carlo Tree Search (MCTS) to efficiently construct a tree of potential questions represented by nodes.
+The question generation process employs a hierarchical approach that combines LLMs with MCTS to efficiently construct a tree of potential questions represented by nodes.
 In the information-seeking phase, at each turn $t$, the system either traverses or expands the tree to find the most optimal questions to ask, with the goal of maximizing information gain.
-This tree is cached and re-used across all data-points belonging to a given dataset.
-The LLM-based question generator takes two inputs: the current possibility set $\Omega_{t}$ and the ancestral context $\mathcal{A}_{\Omega t}$ which consists of the sequence of question-answer pairs corresponding to all the ancestor nodes of the current node that led to the formation of $\Omega_{t}$.
+This tree is cached and reused across all data-points belonging to a given dataset.
+The root node represents the initial conversational state in the decision tree, containing the complete predefined set of possibilities $\Omega$ in a dataset or specific domain.
+Over each of the question-answer turns, some of these possibilities are eliminated by asking information-seeking questions.
+Starting with a common, fixed superset of possibilities allows for efficiency, as the generated questions (child nodes) can be reused for future samples.
+This can lead to more general or broadly applicable questions in the beginning.
+However, for more relevant questions, we can initialize the root node as a constrained set of possibilities $\Omega_{c}\subseteq\Omega$, conditioned by a user’s problem description.
+Specifically, when a new problem arrives, the LLM is prompted to classify likely and unlikely outcomes in $\Omega$ based on the user’s problem.
+Upon determining $\Omega_{c}$, we search the existing first layer of tree for a root node whose possibility set $\Omega_{root}$ is a superset of the constrained set, i.e., $\Omega_{root}\supseteq\Omega_{c}$.
+If such a root node is found, we initiate the questioning process from that node, leveraging the existing question paths or generating more from there.
+Otherwise, a new root node is added to the tree, initialized with the constrained set $\Omega_{c}$.
+This can lead to a wider tree, as there are multiple possible root nodes, essentially representing different subsets of $\Omega$.
+We report results with both types of initializations.
+The LLM-based question generator takes two inputs at turn $t$: the current possibility set $\Omega_{t}$ and the ancestral context $\mathcal{A}_{\Omega t}$ which consists of the sequence of question-answer pairs corresponding to all the ancestor nodes of the current node that led to the formation of $\Omega_{t}$.
 This context prevents the generation of redundant or previously asked questions, thereby providing a clearer path of deduction.
@@
 For each generated question, $LLM_{gen}$ performs a binary partitioning of the possibility set $\Omega_{v}$ into two disjoint subsets $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$, corresponding to affirmative and negative responses respectively, such that $\Omega_{v}=\Omega_{v}^{A}\cup\Omega_{v}^{N}$.
-The LLM is prompted to generate questions that maximize information gain by creating balanced partitions where:
-$|\Omega_{v}^{A}|-|\Omega_{v}^{N}|\rightarrow 0$ (2)
+The LLM is prompted to generate questions that maximize information gain by creating balanced partitions where: $|\Omega_{v}^{A}|-|\Omega_{v}^{N}|\rightarrow 0$.
 Each question spawns two child nodes $v^{A}$ and $v^{N}$ corresponding to affirmative and negative responses, storing $\Omega_{v}^{A}$ and $\Omega_{v}^{N}$ respectively.
-For example, if the possibility set is $\Omega_{v}=\{\text{flu},\text{pneumonia},\text{enteritis},\text{asthma}\}$, the question “Do you have difficulty breathing?” might partition the set as: $\Omega_{v}^{A}=\{\text{pneumonia},\text{asthma}\},\quad\Omega_{v}^{N}=\{\text{flu},\text{enteritis}\}$.
-3.3 Uncertainty-Based Reward
-To quantify the effectiveness of each question in reducing uncertainty, entropy-based metrics are employed Hu et al.
-(2024).
+For example, if the possibility set is $\Omega_{v}=\{\text{flu},\text{pneumonia},\text{enteritis},\text{asthma}\}$, the question “Do you have difficulty breathing?" will partition the set as: $\Omega_{v}^{A}=\{\text{pneumonia},\text{asthma}\},\quad\Omega_{v}^{N}=\{\text{flu},\text{enteritis}\}$.
+3.3 Information Gain
+To quantify the effectiveness of each question in reducing uncertainty, entropy-based metrics are employed ([7]).
 Following the formulation of reward structure in UoT, let $p_{v}^{A}=|\Omega_{v}^{A}|/|\Omega_{v}|$ and $p_{v}^{N}=1-p_{v}^{A}$.
@@
 To normalize and sharpen rewards, a scaling parameter $\lambda>0$ is employed.
-Formally, $R_{IG}(v)$ is defined as:
-$R_{IG}(v)=\frac{-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}}{1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|}.$ (4)
-The expected reward of asking a question, $R_{e}(v)$, is obtained by recursively adding the immediate reward $R_{IG}(v)$ (i.e., information gain at the current node) and the expected rewards of its child nodes.
+Formally,
+$R_{IG}(v)=\frac{-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N}}{1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|}$ (3)
+The expected reward of asking a question, $R_{e}(v)$, is obtained by recursively adding the immediate reward $R_{IG}(v)$ (information gain at the current node) and the expected rewards of its child nodes.
 The total information gained until node $v$ of an interaction trajectory is denoted by the accumulated reward, $R_{a}(v)$, which is obtained by starting at the root and propagating down to the node $v$.
+$R_{a}(v):=R_{IG}(v)+\begin{cases}0&v\text{~{}is~{}root},\\ R_{a}(\textit{parent}(v))&\text{otherwise.}\end{cases}$ (4)
 […]
-$R_{a}(v):=R_{IG}(v)+\begin{cases}0&v\text{~{}is~{}root},\\ R_{a}(\textit{parent}(v))&\text{otherwise.}\end{cases}$ (6)
 where $v^{A}$ and $v^{N}$ are the child nodes of $v$ corresponding to affirmative and negative responses respectively.
-For each non-terminal response-specific node $v^{\alpha}$, $\alpha\in\{A,N\}$, the expected reward is computed as the average of the expected rewards of its child nodes, denoted by $\textit{children}(v^{\alpha})$.
-Formally:
+For each non-terminal response-specific node $v^{\alpha}$, $\alpha\in\{A,N\}$, the expected reward $R_{e}(v^{\alpha})$ is the average expected reward of its child nodes, denoted as $\textit{children}(v^{\alpha})$.
+Formally,
 $R_{e}(v^{\alpha})=\frac{1}{|\textit{children}(v^{\alpha})|}\sum_{v^{\prime}\in\textit{ children}(v^{\alpha})}R_{e}(v^{\prime})$ (6)
-3.4 Monte Carlo Tree Search
+3.4 Monte Carlo Tree Search (MCTS)
 The decision tree of questions is constructed iteratively when MCTS is executed over $K$ iterations at each decision-making step across samples. Each iteration includes four phases:
@@
 $UCT(v)=\frac{R_{\text{total}}(v)}{N_{v}}+C\sqrt{\frac{\ln N_{p}}{N_{v}}}$ (7)
-where $R_{a}(v)$ is the accumulated reward, $N_{v}$ is the visit count of node $v$, $N_{p}$ is its parent node’s visit count, and $C$ is the exploration constant.
+where $R_{total}(v)$ refers to the cumulative reward, $N_{v}$ is the visit count of node $v$, $N_{p}$ is its parent node’s visit count, and $C$ is the exploration constant.
+Note that $R_{total}(v)$ is initialized to 0 in the beginning and updated by adding expected reward $R_{e}(v)$ over time.
 2.
@@
 3.5.1 Cluster Assignment and Creation
-When a new sample arrives, its text embedding is created and compared with existing cluster medoids to determine its assignment.
+When a new sample (user) arrives, we generate a text embedding of their problem description.
+This embedding is then compared with existing cluster medoids to determine its assignment.
 A similarity threshold $\tau$ is used to decide whether to assign the embedding to an existing cluster or create a new one.
@@
 If no existing cluster satisfies this condition, a new cluster is created with $\mathbf{e}$ as its initial medoid.
-If assigned to an existing cluster, $\mathbf{e}$ is added to that cluster’s embeddings, and the medoid $\mathbf{m}_{k}$ is recomputed as:
+If assigned to an existing cluster, $\mathbf{e}$ is added to that cluster’s set of embeddings $\mathcal{C}_{k}$, and the medoid $\mathbf{m}_{k}$ is recomputed as:
 $\mathbf{m}_{k}=\underset{\mathbf{x}\in\mathcal{C}_{k}}{\text{argmax}}\sum_{\mathbf{y}\in\mathcal{C}_{k}}\frac{\mathbf{x}\cdot\mathbf{y}}{\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}}$ (11)
-where $\mathcal{C}_{k}$ is the set of embeddings in cluster $k$.
 3.5.2 Bonus Rewards for Feedback Propagation
 For each node $v$, we maintain a dictionary $B_{k}(v)$ which maps each cluster $k$ to the bonus reward (initialized as zero) earned on reaching the target successfully.
-Once a target is correctly identified, $B_{k}(v)$ is updated for each node along the path from the current node back to the root.
+Once the target ($\omega$) is correctly identified, $B_{k}(v)$ is updated for each node along the path from the current node back to the root, if $\omega\in\Omega_{v}$.
 $B_{k}(v)$ is a proportion of the node’s cumulative reward and depends on the absolute depth of the node, given by $d_{v}$ controlled by an exponential decay as we move upward in the tree.
@@
 $B_{k}(v)\leftarrow B_{k}(v)+\beta\cdot R_{\text{total}}(v)\cdot\gamma^{d_{v}}$ (12)
-where $\beta$ is a task-specific scaling factor (denoting proportion of total reward), and $\gamma\in(0,1)$ is a decay factor that controls the influence of the bonus, keeping it higher for early-stage questions which have proven effective to narrow down the large possibility set, and lower for questions near the terminal nodes which are often more specific to each case.
+where $\beta$ is a task-specific scaling factor (denoting proportion of total reward), and $\gamma\in(0,1)$ is a decay factor that controls the influence of the bonus.
+$B_{k}(v)$ is a scalar value that is higher for early-stage questions which have proven effective to narrow down the large possibility set, and lower for later-stage questions near the terminal nodes which are often more specific to each unique case.
 To incorporate these bonus rewards into decision-making, we modify the UCT formula used in the selection step to handle similar cases while maintaining exploration capabilities.
-We introduce a cluster-specific bonus term to bias selection toward nodes associated with successful outcomes for similar data points.
+We introduce a cluster-specific bonus term to bias selection toward question nodes that led to successful outcomes for similar data points in the past.
 The modified UCT formula becomes:
 $UCT_{\text{feedback}}(v,k)=\frac{R_{\text{total}}(v)}{N_{v}}+C\sqrt{\frac{\ln N_{p}}{N_{v}}}+B_{k}(v)$ (13)
-This modification prioritizes nodes from clusters associated with successful outcomes when processing samples similar to their medoids.
-Algorithm 1 shows the end-to-end summary of the proposed framework.
+4 Experiments
+In this section, we provide an overview of the datasets, baselines, evaluation metrics, and implementation details. Algorithm 1 shows the pseudocode of the proposed framework.
+4.1 Datasets
+We demonstrate the effectiveness of our approach by focusing on three diverse domains: Medical Diagnosis (MD), Troubleshooting (TS), and 20 Questions.
+We use the following datasets pre-processed by [7].
+In Medical Diagnosis, a patient initially reports a brief description of their symptoms, based on which the doctor asks questions to diagnose the disease.
+The maximum number of turns $T$ was limited to 6 in the experiments.
+Two datasets were used.
+The DX dataset [19] contains 104 doctor-patient dialogues and five diseases in its test set.
+The MedDG dataset, which originally included over 17,000 conversations across 15 disease types, was refined by removing inconsistent samples.
+We used 454 high-quality samples for evaluation.
+Open-ended responses are allowed in MedDG to test the system’s generalization capabilities in less constrained scenarios.
+Both datasets limit interactions to 6 turns.
+In the Troubleshooting domain, customer support technicians interact with users to identify faults in systems such as cars or electronic devices.
+We use the FloDial dataset [13], containing 153 dialogues across 153 unique fault types.
+Maximum number of turns $T$ was limited to 20.
+In the 20 Questions domain, the task involves identifying a target item by asking up to 20 yes-or-no questions.
+The Common dataset includes 111 items spanning categories such as animals, places, food, and objects, and the Things dataset [6], was filtered to include 200 items.
 Algorithm 1 MISQ-HF
@@
 4:    $\mathbf{e}_{s}\leftarrow\text{Embedding}[\text{{description}}(s)]$
-5:     $\mathcal{C}_{k}\leftarrow\text{AssignCluster}(\mathbf{e}_{s},\mathcal{C},\tau)$triangleright Use Eq. 11-12
+5:    $\mathcal{C}_{k}\leftarrow\text{AssignCluster}(\mathbf{e}_{s},\mathcal{C},\tau)$triangleright Use Eq. 10-11
 6:    Initialize $\Omega_{v}=\Omega$ at current root node $v$
@@
 8:    while $t<T$ and $\textit{target}(s)\textit{ undetected}$ do
-9:         if $t<\delta\cdot T$ and $|\Omega_{v}\geq 2|$ then
-10:              $v^{*}\leftarrow\text{MCTS}(v,K,\mathcal{C}_{k})$triangleright Use Eq. 4-9, 14
+9:         if $t<\delta*T$ and $|\Omega_{v}\geq 2|$ then
+10:             $v^{*}\leftarrow\text{MCTS}(v,K,\mathcal{C}_{k})$triangleright Use Eq. 3-8, 13
 11:             $q_{t}\leftarrow\textit{question}(v^{*})$triangleright Use Eq. 9
@@
 28:end for
-Model Method $\Omega$-aware DX MedDG FloDial
+4.2 Baselines
+We evaluate two primary baselines: Direct Prompting (DP) and the Uncertainty of Thoughts (UoT) framework [7].
+DP directly queries the LLM to generate the next question without structured planning, serving as a minimal-effort benchmark to highlight improvements achieved by strategic decision-making.
+UoT uses tree-based planning to expand question paths that maximize information gain.
+While effective in reducing uncertainty, its exhaustive tree exploration can be expensive.
+This baseline assesses the efficiency improvement of the proposed MISQ-HF approach, which selectively expands promising branches.
+Comparisons with Chain-of-Thought (CoT) [18] and Tree-of-Thoughts (ToT) [20] are omitted, as prior work [7] shows UoT consistently outperforms these methods.
+4.3 Evaluation
+We use three key metrics to evaluate our system’s effectiveness and efficiency.
+First, Success Rate (SR) measures the percentage of cases where the system correctly identifies the target within the maximum allowed turns.
+Second, Mean Conversational length in Successful Cases (MSC) to track the number of turns required.
+Finally, we introduce a novel metric to track Question Generation Calls (QGC), defined as the number of LLM prompting calls required for question generation during planning.
+4.4 Experimental Setup
+Models.
+We employed three different LLMs of varying sizes as the Questioner.
+Llama 3.3 70B Instruct [4] and Mixtral 8*7B Instruct [9] were accessed via the AWS Bedrock [1].
+GPT-4o was accessed via API from OpenAI [11].
+The user (Answerer) was simulated by Llama 3.3 70B Instruct in all tasks.
+We prompted the model with ground truth details (e.g., the patient’s disease or fault description) and maintained a separate conversation history.
+The temperature was set to 0.
+The ratio of turns was defined by $\delta=0.6$.
+MCTS implementation.
+We set the number of iterations $K=10$ and exploration constant $C=0.2$.
+Maximum simulation depth $d_{s}$ was set to 3 to balance computational efficiency with search effectiveness.
+For each $\Omega_{v}$, the LLM was prompted to generate $m=3$ potential questions to maintain diversity.
+For the reward calculation in $R_{IG}(v)$, the scaling parameter $\lambda$ was set to 0.4.
+Experiments were run on an 8-core CPU with 16 GB RAM.
+Feedback mechanism.
+We used a decay factor $\gamma=0.9$ for the bonus rewards.
+The cluster similarity threshold $\tau$ was set to 0.9 in terms of cosine similarity, and the bonus scaling factor $\beta$ was set to 0.2 for all tasks.
+Problem descriptions were embedded using DistilBERT [14] for the troubleshooting domain, and Clinical-BERT [17] for medical diagnosis.
+To optimize computational efficiency, we cache the decision tree, allowing reuse across all samples within each dataset.
+The system tracks Mean QGC through a thread-safe counter.
+Table 1: Results on MD and TS domains in a Closed Set scenario.
+$\Omega$-aware denotes whether the Questioner Model was informed about the possibility space $\Omega$ at the beginning of the conversation.
+Model Method $\Omega$-aware MD: DX MD: MedDG TS: FloDial
 […]
@@
 UoT 82.69 1.18 0.17 88.79 2.03 1.81 59.47 8.14 41.86
-MISQ 87.50 1.97 0.03 89.20 3.46 0.60 74.50 10.15 4.10
-MISQ-HF 99.03 2.19 0.05 90.30 3.42 0.41 72.54 10.36 2.94
-Table 1: Results on Medical Diagnosis and Troubleshooting Domain in a Closed Set scenario.
-$\Omega$-aware denotes whether the Questioner Model was informed about the possibility space $\Omega$ once at the beginning of the conversation.
-4 Experiments
-In this section, we provide an overview of our experiments, including descriptions of the datasets, baselines, evaluation metrics, and implementation details.
-4.1 Datasets
-We aim to demonstrate the effectiveness and versatility of our approach by focusing on three diverse domains: Medical Diagnosis, Troubleshooting, and 20 Questions.
-We use the following datasets pre-processed by Hu et al.
-(2024).
-In Medical Diagnosis, a patient initially reports a brief description of their symptoms, based on which the doctor asks questions to diagnose the disease.
-The maximum number of turns $T$ was limited to 6 in the experiments.
-Two datasets were used.
-The DX dataset by Xu et al.
-(2019) contains 104 doctor-patient dialogues and five diseases in its test set.
-The MedDG dataset, which originally included over 17,000 conversations across 15 disease types, was refined by removing inconsistent samples.
-We used 454 high-quality samples for evaluation.
-Open-ended responses are allowed in MedDG to test the system’s generalization capabilities in less constrained scenarios.
-Both datasets limit interactions to 6 turns.
-In the Troubleshooting domain, customer support technicians interact with users to identify faults in systems such as cars or electronic devices.
-We use the FloDial dataset by Raghu et al.
-(2021), containing 153 dialogues across 153 unique fault types.
-Maximum number of turns $T$ was limited to 20.
-In the 20 Questions domain, the task involves identifying a target item by asking up to 20 yes-or-no questions.
-The Common dataset includes 111 items spanning categories such as animals, places, food, and objects, and the Things dataset Hebart et al.
-(2019), was filtered to include 200 items.
-4.2 Baselines
-The proposed system is compared with two primary baselines: Direct Prompting (DP) and the Uncertainty of Thoughts (UoT) framework Hu et al.
-(2024).
-DP involves querying the LLM directly to generate the next question without leveraging any structured planning mechanism.
-This approach evaluates the basic capability of the LLM to ask effective questions without additional search strategies.
-It serves as a minimal-effort benchmark, highlighting the improvements achieved by strategic decision-making approaches.
-UoT implements a tree-based planning method that performs an exhaustive expansion of all possible question paths to maximize information gain.
-While UoT demonstrates strong performance in reducing uncertainty and improving task success rates, its reliance on full tree exploration up to a predefined depth can lead to significant computational overhead, particularly within large possibility spaces.
-This baseline is included to measure the improvement in efficiency achieved by the proposed MISQ-HF approach, which selectively expands promising branches rather than exploring all possibilities exhaustively.
-Comparisons with methods such as Chain-of-Thought (CoT) Wei et al.
-(2022) and Tree-of-Thoughts (ToT) Yao et al.
-(2023) are omitted, as prior work by Hu et al.
-(2024) has demonstrated that UoT consistently outperforms these techniques in both efficiency and effectiveness.
-4.3 Evaluation
-We use three key metrics to assess both the effectiveness and efficiency of the question-asking system.
-First, we measure Success Rate (SR), defined as the percentage of cases where the system correctly identifies the target item within the maximum allowed turns.
-Second, we track Mean Conversation Length in Successful Cases (MSC), which represents the average number of turns needed to reach the correct answer (target) in successful interactions.
-Finally, we introduce a novel metric to track Question Generation Calls (QGC), defined as the number of LLM prompting calls required for question generation.
-Lower values indicate better computational efficiency.
-This metric directly measures the system’s ability to minimize expensive LLM operations while maintaining performance.
-Together, these metrics provide a comprehensive view of the system’s performance, balancing success rate, conversation efficiency, and computational overhead.
-4.4 Experimental Setup
-Models.
-We employed three different LLMs of varying sizes as the Questioner.
-Llama 3.3 70B Instruct Grattafiori et al.
-(2024) and Mixtral 8*7B Instruct Jiang et al.
-(2024) were accessed via the AWS Bedrock API (2023).
-GPT-4o was accessed via API provided by OpenAI et al.
-(2024).
-The user (Answerer) was simulated by Llama 3.3 70B Instruct in all tasks.
-We prompted the model with ground truth details (e.g., the patient’s disease or fault description) and maintained a separate conversation history.
-The temperature was set to 0.
-MCTS implementation.
-We set the number of iterations $K=10$ and exploration constant $C=0.2$.
-The maximum simulation depth $d_{s}$ was set to 3 to balance computational efficiency with search effectiveness.
-For each $\Omega_{v}$, the question generator creates $m=3$ potential questions to maintain diversity.
-For the reward calculation in $R_{IG}(v)$, the scaling parameter $\lambda$ was set to 0.4, for both MCTS and UoT.
-Feedback mechanism.
-We used a decay factor $\gamma=0.9$ for the bonus rewards, ensuring early-stage questions receive higher weights.
-The cluster similarity threshold $\tau$ was set to 0.9 in terms of cosine similarity, and the bonus scaling factor $\beta$ was set to 0.2 for all tasks.
-Problem descriptions were embedded using DistilBERT Sanh et al.
-(2019) for the troubleshooting domain, and Clinical-BERT Wang et al.
-(2023) for medical diagnosis.
-To optimize computational efficiency, we cache the decision tree, allowing reuse across all samples within each dataset.
-The system tracks Mean QGC through a thread-safe counter.
+MISQ 87.50 1.97 0.05 89.20 3.46 0.60 74.50 10.15 4.10
+MISQ-HF 99.03 2.19 0.03 90.30 3.42 0.41 72.54 10.36 2.94
+Table 2: Results on MD and TS Domain in a Closed Set scenario, when the root node is initialized with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$.
+Leads to improvement in SR and MSC.
+Model $\Omega$-aware MD: DX MD: MedDG TS: FloDial
+[…]
+Llama 3.3 70B Instruct $\times$ 85.57 0.93 0.20 80.61 2.81 6.13 43.13 6.83 28.26
+98.07 0.96 0.21 89.20 2.67 4.09 72.54 8.25 23.35
+Mixtral 8*7B Instruct 90.38 1.48 0.24 89.87 2.27 4.69 59.47 7.18 30.30
+GPT-4o 99.03 0.54 0.05 92.95 2.15 3.69 78.43 7.18 13.47
 5 Results and Discussion
-Table 1 shows our results in the Closed Set scenario.
+Table 1 shows results in the Closed Set scenario.
 Our approach requires fewer QGC while maintaining higher SR compared to baseline methods, indicating both improved effectiveness and computational efficiency.
 Notably, the $\Omega$-aware Questioner consistently performs better as compared to when unaware.
-To demonstrate the importance of the feedback mechanism, we also report results using MISQ, our framework without the hierarchical feedback component, to rationalize the design choice and support an ablation study.
-In the Medical Diagnosis domain, MISQ-HF achieves superior performance across both datasets, with an average reduction of 6.77 times in Mean QGC across all three LLMs.
-Using Llama 3.3 70B as a Questioner, MISQ-HF required only 0.04 Mean QGC compared to UoT’s 0.77 on DX, showing a 19.25 times reduction.
+To demonstrate the importance of the feedback mechanism, we also report the results using MISQ, our framework without the hierarchical feedback component, to rationalize the design choice and support an ablation study.
+Figure 2: Gain in SR when using MISQ-HF initialized with a constrained set.
+Across both datasets in the MD domain, MISQ-HF achieves superior performance, with an overall average reduction of 6.77 times in Mean QGC over UoT.
+Notably, using Llama 3.3 70B, MISQ-HF required only 0.04 Mean QGC compared to UoT’s 0.77 on DX, showing a 19.25 times reduction.
 Similarly, a reduction of 8.97 times is observed on MedDG.
@@
 While MSC is marginally higher for MISQ-HF compared to UoT, this trade-off can be overseen by the higher success rate of diagnosis, which directly impacts patient outcomes.
-In the Troubleshooting domain using the FloDial dataset, MISQ-HF continues to excel with higher SR, surpassing UoT by 20.68% improvement on average across the three LLMs.
+In the TS domain using FloDial dataset, MISQ-HF continues to excel with higher SR, surpassing UoT by 20.68% improvement on average across the three LLMs.
 Our framework also achieves notable computational efficiency as compared to UoT, with an average of 18.63 times reduction in Mean QGC.
 While troubleshooting conversations typically require more turns due to fault complexity, MISQ-HF maintains a similar MSC (a difference of less than 3 turns on average) while delivering improved performance.
-With GPT-4o, MISQ-HF achieved an SR of 72.54% with 27 clusters, the largest consisting of 31 samples.
-Table 2 shows performance on the 20 Questions domain, evaluated using the Common and Things datasets.
-Due to a lack of problem description or initial hints about the target, we do not evaluate MISQ-HF on this domain.
-For both datasets, MISQ consistently outperforms UoT and DP in all three metrics.
-These results demonstrate the scalability of the proposed approach to larger possibility spaces.
+Table 2 reports the performance when we initialized the root as a constrained set of possibilities.
+This generates more relevant questions early on, based on the user’s problem description.
+Figure 2 shows the notable gain in SR using GPT-4o.
+On average, across the three LLMs on MD datasets, we find that SR is improved by 27.52% over DP, 11.91% over UoT and 5.58% over MISQ-HF initialized with full superset.
+Across both domains, an average of $\approx 8\%$ gains were observed using constrained set.
+This comes with a trade-off in QGC, as the tree gets wider and increases the planning requirements.
+Additional analysis, example conversations and prompts are provided in the Appendix.
+Table 3: Results on 20 Questions Data: Closed Set
 Method $\Omega$- aware Common Thing
@@
 MISQ 85.58 8.51 4.86 55.50 9.54 2.19
-Table 2: Results on 20 Questions Datasets in the Closed Set scenario
+Table 3 shows performance on the general information-seeking domain, evaluated using 20-Questions on Common and Things datasets.
+Due to a lack of problem description or initial hints about the target, we cannot evaluate MISQ-HF here.
+For both datasets, MISQ consistently outperforms UoT and DP in all three metrics.
+These results demonstrate the scalability of the proposed approach to larger possibility spaces.
+Results in the Open Set scenario are provided in Table 5 of Appendix.
+Our extensive evaluation shows that the MISQ-HF outperforms other baselines on datasets with different sizes of possibility spaces.
+The hierarchical feedback mechanism enables the system to learn from successful questioning trajectories, especially in domains like troubleshooting and medical diagnosis, where similar cases often recur.
+Performance without hierarchical feedback (MISQ) further emphasizes the benefits of this approach.
+Compared to exhaustive methods like UoT (Table 1), MISQ-HF achieves $\approx$10x reduction in Mean QGC while delivering $\approx 12\%$ improvement in SR across these domains.
+Limitations.
+The system currently does not incorporate a mechanism to learn from mistakes in failure cases, which could potentially refine its decision-making process and adapt to edge cases more effectively.
+Penalizing redundant or suboptimal quality of questions also requires careful design of reward function and remains an open challenge.
+Incorporating confidence metrics to better quantify uncertainty in risk-sensitive domains is another potential improvement.
+Addressing these limitations presents a promising avenue for future research.
+6 Conclusion
+In this paper, we introduced Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback (MISQ-HF), a novel framework that addresses fundamental limitations in adaptive question-asking for goal-oriented conversational systems.
+Our approach makes three key contributions to the literature on conversational planning: (1) a principled inference-time planning procedure for selecting optimal questions via selective tree expansion, (2) a hierarchical feedback mechanism that incorporates historical performance signals in the UCT formulation to modulate exploration-exploitation trade-offs, and (3) an effective and efficient method that reduces the number of LLM calls made during the planning phase.
+Empirical validation across domains such as medical diagnosis, technical troubleshooting, and general information seeking demonstrates substantial improvements in task performance, while reducing and resource utilization by an order of magnitude.
+The observed gains in constrained possibility spaces further suggest promising avenues for integration with domain-specific knowledge structures.
+The broader implications of our work extend beyond immediate performance metrics to address fundamental challenges in human-machine collaborative problem-solving.
+By combining structured planning with adaptation mechanisms derived from historical interactions, we establish a foundation for conversational AI systems capable of reasoning under uncertainty.
+Future directions include extending MISQ-HF to multi-agent settings, integrating explainability into planning, and investigating theoretical convergence under varying domain characteristics.
+We posit that MISQ-HF represents a significant step toward conversational systems that can engage in truly adaptive, efficient, and goal-directed information acquisition—a critical capability for next-generation AI assistants.
+Appendix
+Appendix A Improvement in Success Rate with $\Omega_{c}$
+Figure 3 demonstrates the % Change reported in Table 4.
+We find that an additional 8% gain in success rate is observed on average when we start with a constrained set of possibilities as compared to full superset of possibilities using MISQ-HF.
+Even larger gains are observed as compared to baselines UoT and DP.
+Figure 3: Improvement in Success Rate on MD and TS Domain in a Closed Set scenario, when initializing the root node with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$.
+Table 4: Improvement in Success Rate on MD and TS Domain in a Closed Set scenario, when using MISQ-HF-c, that is, initialization of root node with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$.
+MISQ-HF-s denotes initialization from full superset at the root node.
+Model Dataset DP UoT MISQ-HF-s MISQ-HF-c
+% Change
+over DP
+% Change
+over UoT
+% Change over
+MISQ-HF-s
+GPT-4o DX 73.07 82.69 99.03 99.03 35.52 19.76 0.00
+MedDG 81.27 88.79 90.3 92.95 14.37 4.68 2.93
+FloDial 43.79 59.47 72.54 78.43 79.10 31.88 8.12
+Llama 3.3 70B Instruct DX 88.46 79.8 98.07 98.07 10.86 22.89 0.00
+MedDG 84.14 89.86 93.39 89.2 6.01 -0.73 -4.48
+FloDial 21.56 60.78 67.97 72.54 236.45 19.35 6.72
+Mixtral 8*7B Instruct DX 50.0 76.92 70.19 90.38 80.76 17.49 28.77
+MedDG 76.43 83.7 84.58 89.87 17.58 7.371 6.25
+FloDial 16.99 39.21 49.01 59.47 250.03 51.67 21.34
+Appendix B Open Set Results
 The results in the Open Set scenario are demonstrated on two datasets where the range of possible outcomes is often unpredictable and varies in size.
@@
 Table 5 shows consistently lower Mean QGC while maintaining similar SR in both domains.
+Table 5: Results in the Open Set scenario. Possibility set $\Omega$ is unknown to the Questioner LLM.
 Method FloDial DX
@@
 MISQ-HF 28.75 6.95 5.10 37.50 2.35 6.32
-Table 3: Results in the Open Set scenario. Possibility set $\Omega$ is unknown to the Questioner LLM.
-[…]
-The frequency of QGC for exhaustive tree search and MCTS differs significantly.
-An LLM call creates $m$ questions, each splitting $\Omega$ into affirmative and negative response sets, thereby giving rise to $2m$ nodes.
-For UoT, the first simulation requires $1+2m+(2m)^{2}+\dots+(2m)^{d_{s}}=\frac{(2m)^{d_{s}+1}-1}{2m-1}$ calls, and $(2m)^{d_{s}}$ calls for subsequent turns if paths are expanded till $d_{s}-1$, where $d_{s}$ is the depth-limit for simulation.
-For MCTS, the first expansion requires 1 call, and simulation for selecting a question at each turn requires $K\cdot d_{avg}$, where $K$ is the number of iterations and $0\leq d_{avg}\leq d_{s}$ because traversing requires 0 LLM calls when the nodes are already expanded.
-Exhaustive search scales exponentially with $d_{s}$, while MCTS scales linearly, highlighting its efficiency for deeper searches in the simulation phase.
-Limitations.
-The system currently does not incorporate mechanisms to learn from mistakes in failure cases, which could potentially refine its decision-making process.
-Explicitly penalizing suboptimal or redundant questions also requires careful design of reward function and remains an open challenge.
-The framework can benefit from incorporating confidence metrics to better quantify uncertainty in risk-sensitive decision making domains.
-Addressing these limitations is a promising direction for future work.
-6 Conclusion
-In this paper, we introduced Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback (MISQ-HF), a novel framework for adaptive question-asking with a goal to optimize conversational efficiency and accuracy.
-By leveraging selective tree expansion and cluster-based feedback, our approach reduces computational overhead while maintaining high success rates across diverse domains such as medical diagnosis, troubleshooting, and general information-seeking tasks.
-We demonstrate the potential of combining structured and efficient planning with historical learning to dynamically adapt question-asking strategies.
-Our contributions aim to refine the decision-making process, enabling conversational AI systems to handle increasingly complex and open-ended scenarios with better planning.
-Acknowledgments
-This work is supported by the National Science Foundation (NSF) award IIS-2336769.
-The authors thank Hafsah Usman, student of the University of Washington, Seattle, for her valuable assistance in running a part of the experiments.
-Appendix A Conversation Prompts for 20 Questions
-A.1 Questioner Prologue
+Appendix C Conversation Prompts for 20 Questions
+C.1 Questioner Prologue
 Let us play the game of 20 questions.
@@
 This is given only once at the beginning of the conversation.
-A.2 Answerer/User Simulator Prologue
+C.2 Answerer/User Simulator Prologue
 Let us play the game of 20 questions.
@@
 Note that you must never reveal X, until I guess it correctly.
-If I guess X correctly in my question, directly respond “You guessed it. X is ‘{target_item}’.” instead of saying yes. Let us begin. Here is my first question.
+If I guess X correctly in my question, directly respond “You guessed it. X is ‘{target_item}’." instead of saying yes. Let us begin. Here is my first question.
 {target_item} the ground truth target item based on which user responds
-A.3 Targeting Prompt
+C.3 Targeting Prompt
 Note that you should guess and ask what X exactly is from now on.
@@
 Target Question: Is X ‘{item}’?
-A.4 Question Generation Prompt
+C.4 Question Generation Prompt
 Here are all the X: {item_set}
@@
 {m} number of potential questions to generate (actions)
-Appendix B Conversation Prompts for Medical Diagnosis
-B.1 Questioner Prologue
+Appendix D Conversation Prompts for Medical Diagnosis
+D.1 Questioner Prologue
 You are a doctor, and your patient self-reports that: {problem_description}.
-{inform_set} You should ask your patient questions in English about symptoms which can only be answered by Yes or No, in order to find what disease this patient suffers from.
-Use the ongoing conversation for context to avoid redundant questions.
+{inform_set} You should ask your patient questions in English regarding their symptoms which can only be answered by Yes or No, in order to diagnose what disease this patient suffers.
+Carefully review the patient’s problem and the ongoing conversation.
+Avoid redundant questions.
 Let us begin.
 Ask me the first question.
-B.2 Answerer/User Prologue
+The text “which can only be answered by Yes or No” is omitted in case of free-answer conversations.
+D.2 Answerer/User Prologue
 You are the patient suffering from ‘{target_item}’, and I am the doctor.
 I will ask you up to 6 questions, and you should answer each one truthfully based on your disease, by saying Yes or No.
-Note that you must never reveal the disease until I tell it correctly.
-If I tell the disease correctly in my question, directly respond: “You are right.
-I am experiencing ‘{target_item}’.” Let us begin.
+Note that you must never reveal the disease until I mention or ask about it.
+If I mention your disease in my question or ask about its symptoms, then you must directly respond “You are right.
+I am experiencing ‘{target_item}’." while saying Yes.
+Let us begin.
 Here is my first question.
-B.3 Question Generation Prompt
+In the beginning of the conversation:
+You are a patient suffering from the disease of ‘{target_item}’, and communicating with a doctor. Here is your conversation history with another doctor: ‘{conv_history}
+Remember the conversation above to answer current doctor’s question in English and do not reveal the disease until the doctor correctly mentions or asks about it.
+If the doctor mentions your disease in their question or asks whether you experience item, you must directly respond “You are right.
+I am experiencing ‘{target_item}’."
+D.3 Question Generation Prompt
 You are a doctor. Here are all the possible diseases that the patient may suffer from: {item_set}
@@
 Count of NO: …
-B.4 Targeting Prompt
+D.4 Targeting Prompt
 Note that you should point out and ask what disease the patient suffers from now.
-The patient is probably suffering from one of the diseases below: {item_set_terminal}, or other from the initial set of possibilities.
-Refer to the past conversation regarding the patient’s symptoms.
-Never repeat previously asked questions.
-Respond with 1 new question only, following the format: “Are you experiencing [disease name]?” Ensure that [disease name] was not asked before.
+The patient is likely suffering from one of the diseases below: {item_set_terminal}, or other from the initial set of possibilities.
+Refer the past conversation about patient’s symptoms.
+Never ask redundant questions.
+Respond with 1 new question only, following the format: “Are you experiencing [disease name]?" Ensure that [disease name] was not asked before.
 Target Question: Are you experiencing ‘{item}’?
-B.5 Classify Prompt
+D.5 Classify Prompt
 Here are all diseases that the patient may suffer from: {item_set}
-{problem_description} For each disease under this report, if the patient is possibly affected, put this disease into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
+{problem_description} According to the patient’s symptoms, classify each possible disease given in the above set into ‘YES: …’ if even slightly possible, otherwise to ‘NO: …’ if completely unlikely.
+Only classify the diseases in above given set, do not add anything else.
+Your answer should be like:
 YES: comma-separated, list of disease names, …
 NO: comma-separated, list of disease names, …
-B.6 Open Set Prompts
+D.6 Open Set Prompts
 Initial Prompt:
@@
 Please propose {size} diseases that you think your patient may suffer from.
-Your response should only be a list like: [”disease1”, ”disease2”, …]
+Your response should only be a list like: [“disease1", “disease2", …]
 Renewal Prompt:
@@
 The list of {size} diseases should contain {existing_items}.
-Your response should only be a list like: [”disease1”, ”disease2”, …]
-Appendix C Conversation Prompts for Troubleshooting
-C.1 Questioner Prologue
+Your response should only be a list like: [“disease1", “disease2", …]
+Appendix E Conversation Prompts for Troubleshooting
+E.1 Questioner Prologue
 You are a technician, and your client self-reports that: {problem_description}.
@@
 Ask me the first question.
-C.2 Answerer/User Prologue
-You are the client with a device that has ‘{target_item}’, and I am the technician.
+E.2 Answerer/User Prologue
+You are the client with a device that has ‘{target_item}’ and I am the technician.
 I will ask you up to 20 questions, and you should answer each one truthfully based on the issue of your device, by saying Yes or No.
@@
 If I tell your issue correctly in my question, directly respond: “You are right.
-My device has issues with ‘{target_item}’.” Let us begin.
+My device has issues with ‘{target_item}’." Let us begin.
 Here is my first question.
-C.3 Question Generation Prompt
+E.3 Question Generation Prompt
 You are a technician. Here are all the issues that the client may face: {items_set}
@@
 Then classify the possible issues above based on this question.
-If the answer is ’YES’, put this issue into ’YES: …’, otherwise to ’NO: …’.
+If the answer is ’YES’, put this issue into ‘YES: …’, otherwise to ‘NO: …’.
 Finally calculate how many issues are in YES and NO.
@@
 Count of NO: …
-C.4 Targeting Prompt
+E.4 Targeting Prompt
 Note that you should now point out and ask what issue the client is facing.
-Refer to the past conversation for context and never ask redundant questions.
-Respond with 1 new question only, following the format: “Are you experiencing [issue name]?” Ensure that [issue name] was not asked before.
+The client is likely to be facing one of the issues below: {item_set_terminal}, so you must consider these.
+Refer the past conversation for problem context.
+Respond with 1 new question only, follow the format: “Are you experiencing [issue name]?" Ensure that [issue name] was not asked before.
 Target Question: Are you experiencing ‘{item}’?
-C.5 Classify Prompt
+E.5 Classify Prompt
 Here are all the issues that the client may face: {item_set}
-{problem_description} For each issue under this report, if the client is possibly facing it, put this issue into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
+{problem_description} According to the user’s problem, classify each possible issue given in the above set into ’YES: …’if even slightly possible, otherwise to ’NO: …’if absolutely unlikely.
+Only classify the issues in above set, do not add any other.
+Your answer should be like:
 YES: comma-separated, list of issue names, …
 NO: comma-separated, list of issue names, …
-C.6 Open Set Prompts
+E.6 Open Set Prompts
 Initial Prompt:
@@
 Please propose {size} issues that you think your client may face with.
-Your response should only be a list like: [”issue1”, ”issue2”, …]
+Your response should only be a list like: [“issue1", “issue2", …]
 Renewal Prompt:
@@
 The list of {size} issues should contain {existing_items}.
-Your response should only be a list like: [”issue1”, ”issue2”, …]
-Appendix D Examples
+Your response should only be a list like: [“issue1", “issue2", …]
+Appendix F Examples
 Two sequences of history are maintained respectively for Questioner LLM and User (Answerer LLM).
@@
 At the beginning of the conversation, the Question LLM is given the Questioner Prologue as the first prompt, and the Answerer LLM is given the Answerer Prologue as the first prompt.
-The rest of the conversation history remains the same, except for the switching between the “system” role and the “user” role for each LLM.
-Table 4, 5, 6 show example conversations from three domains.
-S.No. System User
-Problem Description:
-Hello doctor, I have been in this situation for several years.
-Every summer, I often have diarrhea, but it has gotten worse in the past year or two.
-There is no regular pattern, sometimes it happens every two to three days when it’s mild, and two to three times a day when it’s severe.
-$\dots$ I went to a major hospital for examination before, and they found no parasite eggs or fungi, just a slight imbalance in the bacterial colonies, but the medication didn’t work.
-I don’t know what my condition is, I wonder if any doctor can provide some answers.
-Thank you.
-(Male, 28 years old).
-1 Do you experience any symptoms related to the respiratory system, such as coughing or difficulty breathing?
-No, I don’t experience any symptoms related to the respiratory system, such as coughing or difficulty breathing.
-My main issue is with my stomach and bowel movements.
-I have been experiencing diarrhea, stomach aches, and urgent bowel movements, especially during the summer.
-The stool is often loose, sticky, and not well-formed, and it’s difficult to flush.
-I’ve been having this issue for several years, and it’s been getting worse over the past year or two.
-2 Do you experience any abdominal pain or discomfort?
-Yes, I do experience abdominal pain and discomfort.
-It usually starts with my stomach aching more and more, and then I have to go to the bathroom.
-The pain can be quite severe, especially when the symptoms are bad.
-Sometimes, the stool is not very loose, but it’s accompanied by mucus and is not well-formed, which makes it difficult to flush.
-The abdominal pain and discomfort are quite frequent, and it’s been a persistent issue for me.
-3 Do you experience any symptoms related to digestion, such as nausea or vomiting?
-I don’t usually experience nausea or vomiting, but I do have diarrhea, which can be quite loose and sticky at times.
-The stool is often not well-formed and can be difficult to flush.
-The frequency of the diarrhea can vary, sometimes it’s every two to three days when it’s mild, and sometimes it’s two to three times a day when it’s severe.
-The symptoms can be quite unpredictable and irregular, which makes it challenging to manage.
-4 Are you experiencing symptoms of irritable bowel syndrome? You are right. I have Irritable bowel syndrome.
-Table 4: An example conversation from MedDG dataset using MISQ-HF with GPT-4o.
+The rest of the conversation history remains the same, except for the switching between the “system" role and the “user" role for each LLM.
+Table 7, 6, 8 show example conversations from three domains.
+Table 6: An example conversation from FloDial dataset using MISQ-HF with GPT-4o.
 S.No. System User
@@
 6 Are you experiencing a radiator leak? You are right. It’s a Radiator leak.
-Table 5: An example conversation from FloDial dataset using MISQ-HF with GPT-4o.
+Table 7: An example conversation from MedDG dataset using MISQ-HF with GPT-4o.
+S.No. System User
+Problem Description:
+At the age of 28, recently there has been a slight pain on both sides of the stomach, the stomach has been bothering for a few days, the stool is slightly bloody, and the stomach is slightly painful, but it comes in waves.
+1 Do you experience symptoms that worsen after eating fatty foods?
+Actually, I’ve been experiencing a burning sensation in my stomach, and it seems to be triggered by certain foods and activities, including drinking alcohol.
+The pain is usually on both sides of my stomach and comes in waves.
+I’ve also had some bloody stools recently.
+However, I’m not sure if fatty foods specifically make it worse, but I do feel like my stomach is more bothered after drinking or eating certain things.
+2 Do you experience abdominal pain that improves after eating?
+The pain in my stomach is a bit unpredictable, but it doesn’t really seem to improve after eating.
+In fact, sometimes I feel like my stomach is burning, and it can be uncomfortable.
+The pain is usually on both sides of my stomach and comes in waves.
+It’s been bothering me for about 20 days now, and I’ve also had some other symptoms like slightly bloody stools.
+3 Are you experiencing a gastric ulcer?
+No, I’m not sure about that.
+I’ve been having some stomach issues, like pain on both sides of my stomach, and it’s been bothering me for about 20 days.
+I’ve also had some other symptoms like slightly bloody stools and a burning sensation in my stomach.
+But I haven’t been diagnosed with a gastric ulcer.
+I did talk to another doctor who suggested I take some medication, omeprazole, to help with my symptoms.
+4 Are you experiencing gastritis? You are right. I have Gastritis.
+Table 8: An example conversation from Common dataset using MISQ with GPT-4o.
 S.No. System User
@@
 8 Is X a ’Chameleon’? You guessed it. X is Chameleon.
-Table 6: An example conversation from Common dataset using MISQ with GPT-4o.
+Appendix G QGC Efficiency
+The frequency of QGC for exhaustive tree search and MCTS differs significantly.
+An LLM call creates $m$ questions, each splitting $\Omega$ into affirmative and negative response sets, thereby giving rise to $2m$ nodes.
+For UoT, the first simulation requires $1+2m+(2m)^{2}+\dots+(2m)^{d_{s}}=\frac{(2m)^{d_{s}+1}-1}{2m-1}$ calls, and $(2m)^{d_{s}}$ calls for subsequent turns if paths are expanded till $d_{s}-1$, where $d_{s}$ is the depth-limit for simulation.
+For MCTS, the first expansion requires 1 call, and simulation for selecting a question at each turn requires $K\cdot d_{avg}$, where $K$ is the number of iterations and $0\leq d_{avg}\leq d_{s}$ because traversing requires zero LLM calls when the nodes are already expanded.
+Exhaustive search scales exponentially with $d_{s}$, while MCTS scales linearly, highlighting its efficiency for deeper searches in the simulation phase.
+Appendix H Broader Impacts
+The proposed framework enhances goal-oriented conversational AI systems in domains like medical diagnosis and troubleshooting.
+By optimizing question-asking strategies, it improves decision-making success rates with reduced computational overhead.
+This can lead to more efficient diagnostic systems and customer support interactions, potentially benefiting both users and service providers.
+While general concerns about privacy and trustworthiness in LLMs persist, they are not specifically pertinent to the proposed system in this work.
+However, we acknowledge that use of LLMs as a core component of a system inherently brings challenges related to privacy and model biases, which must be carefully managed to ensure responsible deployment and maintain user trust.